### v2

Abstract
Electronic Health Records (EHRs) offer rich potential for clinical prediction, yet their inherent complexity and heterogeneity pose significant challenges for traditional machine learning approaches. Domain-specific EHR foundation models trained on large collections of unlabeled EHR data have demonstrated promising improvements in predictive accuracy and generalization; however, their training is constrained by limited access to diverse, high-quality datasets and inconsistencies in coding standards and healthcare practices. In this study, we explore the possibility of using general-purpose Large Language Models (LLMs) based embedding methods as EHR encoders. By serializing patient records into structured Markdown text, transforming codes into human-readable descriptors, we leverage the extensive generalization capabilities of LLMs pretrained on vast public corpora, thereby bypassing the need for proprietary medical datasets. We systematically evaluate two state-of-the-art LLM-embedding models, GTE-Qwen2-7B-Instruct and LLM2Vec-Llama3.1-8B-Instruct, across 15 diverse clinical prediction tasks from the EHRSHOT benchmark, comparing their performance to an EHR-specific foundation model, CLIMBR-T-Base, and traditional machine learning baselines. Our results demonstrate that LLM-based embeddings frequently match or exceed the performance of specialized models, even in few-shot settings, and that their effectiveness scales with the size of the underlying LLM and the available context window. Overall, our findings demonstrate that repurposing LLMs for EHR encoding offers a scalable and effective approach for clinical prediction, capable of overcoming the limitations of traditional EHR modeling and facilitating more interoperable and generalizable healthcare applications.
1 Introduction
EHRs have become a widespread technology in modern healthcare, providing a comprehensive, longitudinal view of a patient’s health status [1]. Machine learning methods can leverage this rich data to perform risk stratification and support clinical decision making [2, 3, 4]. In recent years, researchers have explored a variety of prediction tasks based on EHRs, including hospital readmission [5, 6], length of hospital stay [6], sepsis onset detection [7, 8], mortality prediction [6, 9], discharge diagnoses [6], and heart failure outcomes [10]. The overarching objective is to harness existing EHR data through machine learning to enhance clinical outcomes and reduce healthcare costs.
However, machine learning on EHR data poses significant challenges due to its inherent complexity. EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11]. As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based methods [6, 12, 13]. To mitigate these issues, recent approaches have employed large-scale foundation models that are pre-trained on unlabeled EHR data using unsupervised learning [14]. Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16]. Treating EHR data as sequences of medical codes, enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13]. However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data. Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural-language tasks [20]. This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications. Consequently, LLMs have demonstrated strong performance in extracting medical concepts [21], summarizing medical texts [22], and predicting medical outcomes [23] even in low-resource settings. However, most modern LLMs, such as GPT [24] or Llama [25], utilize a decoder-only transformer architecture, which complicates the generation of robust text representations. To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [26, 27, 28, 29]. Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
In this study, we systematically evaluate whether general-purpose LLM-embedding models can effectively encode EHR records for 15 distinct clinical prediction tasks [30] (see Fig. 1). To this end, we first transform the EHR data into a concise text representation capturing the most relevant patient information at prediction time. We then use two LLM-embedding models, GTE-Qwen2-7B-Instruct [31, 29] and LLM2Vec-Llama3.1-8B-Instruct [26, 32], to generate EHR representations. The resulting embeddings serve as inputs to a logistic regression classifier, which is trained separately for each prediction task. We focus on evaluating the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness.
2 Results
2.1 EHR Database and Prediction Tasks
We used the EHRSHOT database containing adult patients from the Stanford Health Care and Lucile Packard Children’s Hospital from 1990 to February 8th, 2023 [30]. This dataset includes comprehensive EHR timelines for 6.739 patients, covering 921.499 visits, and 41.661.637 clinical events (Table 1). The database is part of a rigorous EHR benchmark containing 15 clinical prediction tasks grouped into four task groups. It also provides canonical dataset splits and provides publicly available code. Table 2 presents a detailed breakdown of the task groups, individual tasks, and the corresponding number of labels used in our experiments. Notably, a single patient may contribute multiple labels for a given prediction task, as relevant clinical events can recur over time.
| Attribute | Train | Validation | Test | Total |
| Num Events | 15,511,472 (37.2) | 13,005,205 (31.2) | 13,144,960 (31.6) | 41,661,637 (100.0) |
| Num Visits | 339,504 (36.8) | 300,325 (32.6) | 281,670 (30.6) | 921,499 (100.0) |
| Num Patients | 2,295 (34.1) | 2,232 (33.1) | 2,212 (32.8) | 6,739 (100.0) |
| Num Female | 1,173 (34.1) | 1,142 (33.2) | 1,126 (32.7) | 3,441 (100.0) |
| Num Male | 1,122 (34.0) | 1,090 (33.1) | 1,086 (32.9) | 3,298 (100.0) |
| Age, mean SD | 59.6 17.8 | 58.8 18.0 | 59.5 18.0 | 59.3 17.9 |
| American Indian | 14 (56.0) | 7 (28.0) | 4 (16.0) | 25 (100.0) |
| Asian | 356 (34.1) | 347 (33.3) | 340 (32.6) | 1,043 (100.0) |
| Black | 98 (32.9) | 105 (35.2) | 95 (31.9) | 298 (100.0) |
| Pacific Islander | 23 (31.1) | 21 (28.4) | 30 (40.5) | 74 (100.0) |
| Unknown | 518 (33.1) | 530 (33.9) | 515 (32.9) | 1,563 (100.0) |
| White | 1,286 (34.4) | 1,222 (32.7) | 1,228 (32.9) | 3,736 (100.0) |
| Hispanic | 374 (36.0) | 342 (32.9) | 322 (31.0) | 1,038 (100.0) |
| Non-Hispanic | 1,921 (33.7) | 1,890 (33.2) | 1,890 (33.2) | 5,701 (100.0) |
| Attribute | Train Labels (Positive) | Valid Labels (Positive) | Test Labels (Positive) | Total Labels (Positive) |
| Operation Outcomes | ||||
| Long Length of Stay | 2,569 (681) | 2,231 (534) | 2,195 (552) | 6,995 (1,767) |
| 30-day Readmission | 2,609 (370) | 2,207 (281) | 2,189 (260) | 7,005 (911) |
| ICU Transfer | 2,402 (113) | 2,052 (92) | 2,037 (85) | 6,491 (290) |
| Anticipating Lab Test Results | ||||
| Thrombocytopenia | 68,776 (9,774) | 54,504 (6,962) | 56,338 (7,960) | 179,618 (24,696) |
| Hyperkalemia | 76,349 (1,215) | 60,168 (886) | 63,653 (948) | 200,170 (3,049) |
| Hypoglycemia | 122,108 (1,065) | 95,488 (858) | 100,568 (783) | 318,164 (2,706) |
| Hyponatremia | 81,336 (20,181) | 64,473 (14,674) | 67,028 (16,003) | 212,837 (50,858) |
| Anemia | 70,501 (9,544) | 56,224 (7,445) | 58,155 (7,636) | 184,880 (24,625) |
| Assignment of New Diagnoses | ||||
| Hypertension | 1,260 (184) | 1,250 (177) | 1,261 (160) | 3,771 (521) |
| Hyperlipidemia | 1,684 (205) | 1,441 (189) | 1,317 (172) | 4,442 (566) |
| Pancreatic Cancer | 2,576 (155) | 2,215 (53) | 2,220 (56) | 7,011 (264) |
| Celiac | 2,623 (62) | 2,284 (11) | 2,222 (21) | 7,129 (94) |
| Lupus | 2,570 (104) | 2,226 (33) | 2,243 (20) | 7,039 (157) |
| Acute MI | 2,534 (175) | 2,177 (146) | 2,127 (144) | 6,838 (465) |
| Anticipating Chest X-ray Findings | ||||
| Chest X-Ray Findings | 7,481 (4,771) | 9,366 (6,032) | 9,428 (6,400) | 26,275 (17,203) |
2.2 EHR Text Serialization
To utilize an LLM-embedding model for EHR representation, the patient records were encoded into a structured text based on the widely used Markdown format. Due to runtime constraints, the maximum serialization length was limited to 4.096 tokens (approximately 16.000 characters). This constraint informed the serialization strategy, prioritizing the inclusion of recent data to ensure critical medical information was preserved. The serialized record is divided into clearly labeled sections (see example in Fig. 2). All dates were normalized relative to a reference prediction date of January 1st, 2024, explicitly stated at the beginning. This is followed by the patient’s basic demographic information. Approximately 65% of the recorded values were time-series data of Logical Observation Identifiers Names and Codes (LOINC) concepts. To reduce the volume of this data, the serialization focuses on 24 frequently recorded concepts, grouped into three primary categories: Body Metrics, Vital Signs, and Lab Results. For each selected concept, the last three recorded values were included, along with the corresponding units and a classification as low, normal, or high, where applicable. The serialization then summarizes all patient visits, including visit type, time and duration, followed by any event not associated with specific visits. Detailed visit records are then provided in descending chronological order, beginning with the most recent visit. Each visit entry is further categorized into conditions, medications, and procedures for improved clarity and utility. As a result, the text serialization ensures a structured and concise representation of EHR data, facilitating efficient processing by LLM-based models.
2.3 LLM-Embedding Models and Baselines
We focus on two LLM-embedding models as part of our experiments to explore their effectiveness in representing EHR data. The first model, GTE-Qwen2-7B-instruct (GTE-Qwen2-7B), is based on the Qwen2-7B LLM [31] and incorporates bidirectional attention and contrastive learning to enhance embedding tasks [29]. The second model, LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B), is built on the Llama-3.1-8B Instruct architecture [32] and employs similar optimization techniques to improve embeddings [26]. Both models were trained using instructions for the embeddings task. Hence, we added a simple prompt for each task, e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia?” (see Table 8). Our primary goal was to assess how these LLM-based models, trained on publicly available text data, perform in representing EHRs compared to an EHR-specific foundation model. For this, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 30]. For each model, we used a logistic regression classification head trained on the training split, with hyperparameters tuned on the validation split. Additionally, we included a counts-based baseline using a Gradient Boosted Machine (GBM) model, which has demonstrated superior performance over logistic regression for EHR tasks [30].
| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
| LLM-Embedding Models | |||||
| GTE-Qwen2-7B | |||||
| GTE-Qwen2-1.5B | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| LLM2Vec-Llama 2 1.3B | |||||
| LLM-Embedding Model + EHR Foundation Model | |||||
| GTE-Qwen2-7B + CLMBR-T-Base | |||||
| LLM2Vec-Llama-3.1-8B + CLMBR-T-Base | |||||
| Baselines [30] | |||||
| CLMBR-T-Base | |||||
| Counts-based + GBM | |||||
| Encoder Language Models | |||||
| DeBERTaV3 large | |||||
| DeBERTa V3 base | |||||
| BERT large | |||||
| BERT base | |||||
| ClinicalBERT | |||||
2.4 Performance Results on 15 Prediction Tasks
The performance results measured by the Area Under the Receiver Operating Characteristic Curve (AUROC) for all training and validation examples are presented in Table 3. The GTE-Qwen2-7B model demonstrated superior performance over CLIMBR-T-Base in three of the four task categories: operation outcomes, lab test result prediction, and assignment of new diagnoses. However, in predicting chest X-ray findings, the EHR foundation model (CLIMBR-T-Base) outperformed GTE-Qwen2-7B. The LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance compared to GTE-Qwen2-7B, surpassing CLIMBR-T-Base only in the assignment of new diagnoses task. Both LLM-based embedding models clearly outperformed the counts-based baseline with a gradient-boosted machine head, which is traditionally regarded as a strong baseline for EHR tasks [12, 13]. The only exception was in the assignment of new diagnoses, where the counts baseline marginally outperformed GTE-Qwen2-7B. We also combined the CLMBR-T-Base embedding with the LLM representations by simple concatenation to test whether they encode orthogonal information. The combined embeddings led to a considerable performance boost for both models with an average AUROC performance of 0.801 (0.777 - 0.826) instead of 0.774 (0.749 - 0.798) for GTE-Qwen2-7B and 0.779 (0.753 - 0.805) instead of 0.742 (0.714 - 0.769) for LLM2Vec-Llama-3.1-8B. The smaller GTE-Qwen2-1.5B model only performed slightly worse than GTE-Qwen2-7B with an average performance of 0.757 (0.732 - 0.783) compared to 0.774 (0.749 - 0.798) making it a performant model choice for EHR embeddings. LLM2Vec-Llama-2-1.3B on the other hand performed significantly worse than the larger model which could be due to the older Llama 2 architecture. Fig. 3 illustrates the relationship between model size and average performance across all tasks, suggesting the trend that larger models tend to perform better. Among the evaluated models, CLIMBR-T-Base proved to be the most efficient in terms of performance relative to model size. It is noteworthy, however, that the LLM-based embedding models were not specifically optimized for this EHR data modality, which may partially explain their performance gap.
2.5 Performance Results in Few-Shot Setting
To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples. The experiments followed the EHRSHOT task definitions [30]. Fig. 4 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples (see Fig. 6 for AUPRC results). Both LLM-embedding models demonstrated strong results in the few-shot setting, indicating that their pretraining on general text can be successfully transferred to serialized EHR data. Notably, GTE-Qwen2-7B consistently outperformed CLIMBR-T-Base across all training example sizes for predicting lab test results and assigning new diagnoses. For operational outcomes, a minimum of 32 training examples was required for GTE-Qwen2-7B to surpass the performance of the EHR-specific foundation model. In contrast, LLM2Vec-Llama-3.1-8B showed weaker performance in the few-shot setting. It only outperformed CLIMBR-T-Base in a limited number of scenarios, primarily for the assignment of new diagnoses. Both LLM-embedding models consistently achieved improvements over the counts-based baseline across most configurations, underscoring the impact of their pretraining on general text. Consistent with previous findings [30], the largest performance gains over the counts-based baseline were observed at intermediate training example sizes. As the number of training examples increased, the advantage of pretrained LLM-based models decreased, highlighting the diminishing returns of foundation models as more labeled data becomes available.
2.6 Effect of Different Contexts Sizes
We evaluated the impact of varying context sizes on the performance of the LLM-embedding models by testing GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B with input lengths of 512, 1.024, 2.048, 4.096 (the default), and 8.192 tokens. The objective was to determine whether longer input sequences enhance performance and to assess if the models effectively leverage additional historical data. As summarized in Table 4 and Fig. 5, the two models exhibited distinct behaviors.
| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
| GTE-Qwen2-7B | |||||
| 8.192 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 8.192 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size |
For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4.096 tokens, with only a slight decline observed when using 2.048 or 1.024 tokens. This suggests that the most relevant information is contained within the first 1.024 tokens and that the model can accurately extract these details even when presented within a longer context. In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2.048 tokens and a moderate gain at 1.024 tokens, indicating that the model struggles to effectively handle longer sequences up to 4.096 tokens. Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [33]. Furthermore, extending the context to 8.192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs. This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
2.7 Effect of Chunked Contexts
To investigate whether the models can process the full 4.096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1.024, and 2.048 tokens. For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 5). For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4.096-token input remains effectively used even when segmented. Notably, using 512-token chunks yielded an overall AUROC performance of 0.735, compared to 0.672 when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints. In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes. This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2.048 tokens.
| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
| GTE-Qwen2-7B | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks |
| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
| GTE-Qwen2-7B | |||||
| no instructions | |||||
| no demographics | |||||
| no aggregated | |||||
| no visits (both) | |||||
| no conditions | |||||
| no medications | |||||
| no procedures | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| no instructions | |||||
| no demographics | |||||
| no aggregated | |||||
| no visits (both) | |||||
| no conditions | |||||
| no medications | |||||
| no procedures |
2.8 Ablations of EHR Serialization
To assess the impact of different components of the EHR serialization, we conducted a series of ablation studies for both LLM-embedding models, as summarized in Table 6. The removal of instructions had a notable effect on performance for both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, leading to a decrease from 0.774 (0.749 - 0.798) to 0.725 (0.700 - 0.750) and from 0.742 (0.714 - 0.769) to 0.727 (0.700 - 0.755) in averaged AUROC, respectively. The most significant drop was observed in tasks related to operational outcomes and lab result prediction, suggesting that the instructions played a critical role in guiding the model to focus on relevant clinical information. The inclusion of aggregated information, comprising four body metrics, six vital signs, and 14 lab values, also had a major impact on predictive performance. Removing this information substantially reduced the accuracy of lab test predictions, with performance dropping from 0.867 (0.860 - 0.874) to 0.713 (0.701 - 0.726) for GTE-Qwen2-7B and from 0.777 (0.766 - 0.787) to 0.705 (0.692 - 0.718) for LLM2Vec-Llama-3.1-8B. This highlights the importance of the most recent lab values in predicting future lab results. The effect was particularly pronounced for GTE-Qwen2-7B in predicting conditions such as thrombocytopenia, hyperkalemia, and hyponatremia. However, for hypoglycemia and anemia, GTE-Qwen2-7B performed on par with CLIMBR-T-Base, even though the lab values were still part of the serialized representation (see Fig. 7 and Fig. 8 for task specific performance). Interestingly, the removal of aggregated information led to slight performance improvements in predicting operational outcomes and the assignment of new diagnoses, suggesting that in some cases, a more focused representation may be beneficial. Visit-related information proved to be particularly critical for operational outcome prediction, as removing it led to a performance drop from 0.844 (0.821 - 0.867) to 0.752 (0.728 - 0.776) for GTE-Qwen2-7B and from 0.787 (0.762 - 0.812) to 0.734 (0.709 - 0.759) for LLM2Vec-Llama-3.1-8B. Notably, the removal of visit information resulted in a substantial performance increase for LLM2Vec-Llama-3.1-8B predicting lab values, from 0.777 (0.766 - 0.787) to 0.862 (0.855 - 0.869). This suggests that the model struggles to extract relevant information when presented with a large volume of contextual data, a challenge that GTE-Qwen2-7B handles more effectively. On the other hand, the removal of demographic information, conditions, medications, and procedures had only a minor effect on performance, indicating that these features were either redundant or less relevant for the prediction tasks. While the influence of aggregated information and visit details was clear for predicting operational outcomes and lab results, the ablation studies showed less impact on tasks related to new diagnoses and chest X-ray prediction. This suggests that these latter task groups rely more on a broader set of contextual factors rather than any single component of the serialization.
3 Discussion
Our study demonstrates that general-purpose LLM-embedding models, originally pre-trained on extensive natural language corpora, can be repurposed as robust foundation models for EHR prediction tasks. Specifically, models like GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B not only surpassed a strong counts-based baseline but, in several clinical domains, matched or even exceeded the performance of a dedicated EHR foundation model (CLIMBR-T-Base) [13, 30]. Notably, this is despite CLIMBR-T-Base being trained on data from the same hospital system as the EHRSHOT database, underscoring the strong generalization capabilities of LLM-based embeddings. This is especially significant given the inherent challenges of limited and heterogeneous EHR data. Across 15 diverse clinical tasks, including few-shot scenarios, our findings indicate that the transferable knowledge acquired during large-scale text pretraining enables these models to effectively capture and encode complex clinical information. Furthermore, our analysis shows that performance scales with both model size and architectural improvements, suggesting that future advancements in general-purpose LLMs may further enhance their applicability in healthcare settings. In contrast, traditional EHR-specific models face scaling limitations due to restricted data availability and the necessity for specialized architectures [34]. Collectively, these results highlight that repurposing LLMs as embedding generators offers a powerful, flexible, and scalable alternative for clinical prediction tasks.
Our experiments suggest that several key factors contribute to the success of LLM-embedding models for encoding EHR data. First, providing clear, task-specific instructions within the serialized EHR text was crucial, as it guided the models to focus on the most clinically relevant sections, particularly benefiting tasks that require attention to specific input segments. This design leverages the general-purpose pretraining and further instruction-based fine-tuning of LLM-embedding models [26, 29], enabling them to identify and extract meaningful patterns from heterogeneous and complex EHR inputs. Second, the inclusion of aggregated clinical data, especially semantic LOINC-coded events and detailed visit-related information, was strongly associated with improved performance on operational outcomes and lab test predictions. In contrast, other components of the EHR did not show as clear a relationship with predictive accuracy. The effective handling of context size also plays a significant role. While models like GTE-Qwen2-7B can fully utilize contexts up to 4,096 tokens, our findings show that performance degrades with overly short or excessively long inputs. Short contexts risk omitting recent, crucial clinical encounters, whereas overly long contexts can dilute important token-level details. In cases where models face challenges with longer inputs such as LLM2Vec-Llama-3.1-8B our chunking experiments indicate that segmenting the input and averaging the resulting embeddings can mitigate this issue. Finally, combining LLM embeddings with those from an EHR-specific foundation model (CLMBR-T-Base) led to a considerable boost in performance. This improvement suggests that LLMs capture orthogonal and complementary clinical information. One explanation for this effect could be the limited context size of CLMBR, restricted to 496 medical codes, which prevents it from fully leveraging all information present in our EHR serialization [13]. Alternatively, the boost may stem from the general-purpose pretraining on vast text corpora, which provides LLMs with domain-agnostic knowledge that remains untapped in models trained exclusively on EHR data.
Unlike count-based models and specialized EHR foundation models, LLM-based models are agnostic towards specific coding systems and data formats. Traditional EHR models depend on predefined vocabularies such as SNOMED, LOINC, and ICD, which restricts their applicability across diverse healthcare settings. In contrast, LLMs operate on raw textual representations, enabling them to interpret any code that can be mapped to a human-readable form. This flexibility is particularly valuable given the persistent challenges in EHR data interoperability, stemming from variable coding practices, privacy constraints, and regulatory limitations, which often hinder the aggregation of large, standardized datasets for model pretraining [35]. Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [36]. This capacity mitigates the limitations of count-based models that typically filter out infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [37]. Additionally, the ability of LLMs to process both structured and unstructured data paves the way for integrated, multimodal representations that could further enhance clinical decision support systems [38].
Our findings highlight that LLM-embedding methods merge the strengths of count-based models, specialized EHR models, and the expansive generalization capabilities of large language models. While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are inherently limited by their reliance on predefined vocabularies and constrained training data. In contrast, LLMs can process any textual representation, enabling them to effectively handle rare or unseen codes, although they typically produce unstructured outputs that challenge clinical grounding and calibration [23]. LLM-embedding methods bridge this gap by leveraging the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks. By transforming EHR data into structured text and embedding it using general-purpose LLMs, we enable the application of logistic regression or other conventional machine learning classifiers to make clinically meaningful predictions. This approach preserves the calibration advantages of traditional models while benefiting from the rich contextual understanding of LLMs.
3.1 Limitations
This study has several limitations. First, our approach relies on a subjectively designed EHR serialization that we deemed to capture the most medically relevant information. This design choice may introduce bias when comparing LLM-based embedding models with dedicated EHR foundation models that operate on raw data. In addition, the performance of the LLM-embedding models is sensitive to the specific content and instructions provided in the serialized text, which may limit reproducibility and generalization ability. Although these models achieve competitive predictive performance across diverse clinical tasks, including few-shot scenarios, their substantially larger parameter counts lead to longer computation times and higher resource usage. Moreover, by relying on LLM embedding methods, we must train a downstream classifier from scratch, thereby forgoing the inherent zero-shot or few-shot prompting capabilities of LLMs. Finally, our serialization was limited to 4,096 tokens to manage runtime constraints, potentially omitting valuable long-range historical information, and our evaluation on a single institutional dataset may limit broader applicability across diverse healthcare systems.
3.2 Future work
Future research should explore serialization‐free approaches that allow LLMs to process raw EHR data directly, thereby reducing potential biases introduced by manual text transformation. Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training. It will also be important to develop strategies to extend the effective context window beyond 4.096 tokens to capture more comprehensive patient histories. Moreover, investigating techniques for distilling large LLMs into smaller, more efficient models may enhance their practical applicability in clinical settings. Finally, expanding evaluations to multi-institutional datasets and examining how complementary insights from both domain-specific EHR models and general-purpose LLMs can be synergistically combined will be critical for advancing the development of robust, scalable EHR foundation models.
4 Methods
4.1 EHR Database and Prediction Task
The EHR data utilized in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [30]. We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement. This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8th, 2023. Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table. We used the official ehrshot-benchmark repository111Github repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments enabling to us to build on existing functionalities and to facilitate comparisons with prior methods. The benchmark uses the Framework for Electronic Medical Records (FEMR) 222Github repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data. All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark. The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical predictions tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [30]. Task labels are derived from clinical events, resulting in significant variations in task-specific sample sizes. For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses. The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pre-trained foundation models [14] due to their ability to generalize from limited training data. To this end, the benchmark defines evaluation settings with a constrained number of training and validation examples. Specifically, for k in 1, 2, 4, 8, 12, 16, 24, 32, 48, 64, 128, the benchmark uses k positive and k negative training examples, along with k positive and k negative validation examples, to train and tune supervised classifiers. Testing is always performed on the full set of examples. The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and gradient boosting machines [39]. Performance is reported using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [30]. Macro averages are reported for each task group, and an overall macro average is provided across all groups.
4.2 EHR Text Serialization
To leverage LLM-embedding models for representing EHR records, we serialized the records into textual formats. We had to limit the length of the serializations to 4.096 tokens (approximately 16.000 characters) to carry out all experiments on the available computing infrastructure (see Section 4.6). Two experimental runs were conducted with serializations extending up to 8.192 tokens. The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text. This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted.
To convert the visits and clinical events in the EHRSHOT dataset into text, we leveraged the semantic information embedded in the dataset. Each clinical event was labeled using the format “ontology/code”. EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated. An analysis was performed to evaluate the utilization of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2.968 labels. We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type. We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type). Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontology. Cancer Modifier code contained UICC cancer stages that we parsed manually. For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx. We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful description.
Various approaches exist for serializing structured data, including row-by-row serialization [40], template-based methods [41], or structured data formats like JSON, HTML, and Markdown [42]. We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [43]. To harmonize dates, all timestamps were normalized relative to January 1, 2024, designated as the prediction reference time. Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity. Since 65% of the dataset comprised time-series data encoded via LOINC, which we found imbalanced, we aggregated LOINC-coded events. Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts. To avoid duplicates, we merged synonymous codes (see Table 7). The last three values of each concept are given, and we filtered implausible values. To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 7). Following the aggregated data, a summary of all visits was included to address the potential truncation of older visits. Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable. Finally, a detailed chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).
4.3 Potential Bias of Manually Defining an EHR Text Serialization
Defining an EHR serialization involved subjective decisions, which may have introduced bias. For instance, awareness of the prediction tasks could influence the prioritization of certain data elements, potentially favoring task-relevant information. To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records. Also, due to computational constraints, we evaluated only three variants of our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig. 2), and (3) combining both approaches. For GTE-Qwen2-7B, the chosen serialization performed best, but all variants outperformed CLMBR-T-Base. However, for LLM2Vec-Llama-3.1-8B, only the serialization omitting all values (variant 2) performed slightly better. We selected the current serialization for its simplicity and completeness of medical information. This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
4.4 LLM-Embedding Models and Baselines
In this study, we evaluated two LLM-embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs. These models were selected for their ability to handle the 4.096-token EHR serializations used in our experiments. For comparison, we also tested a smaller variant of both models. As additional baselines we included commonly used encoder-only embedding models with smaller input sizes (512 tokens). To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation. For all LLM-embedding models and smaller language models used in this study, we used the mean pooling of the last layer as the final embedding [26, 44]. The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction. Below is an overview of all models:
GTE-Qwen2-7B
This LLM-embedding model is based on the Qwen2-7B-Instruct LLM [31] using a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3.584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct. It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [29]. This conversion replaces causal attention with bidirectional attention, enabling the model to attend to both left and right contexts for token embedding, akin to BERT. Contrastive learning was applied using a mixture of private datasets to enhance embedding performance. The model also incorporates instructions tailored for embedding tasks, supporting a context size of up to 32.000 tokens.
GTE-Qwen2-1.5B
A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536. It was also trained using the GTE method [29] and supports a context size of up to 32.000 tokens.
LLM2Vec-Llama-3.1-8B
This model is built upon the Llama-3.1-8B-Instruct LLM [32] with a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4.096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [26]. This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks. The finetuning used curated data from the public E5 dataset [45, 46], containing approximately 1.5 million entries. The model supports task-specific instructions and accommodates a context size of up to 128.000 tokens.
LLM2Vec-Llama-2-1.3B
DeBERTa v3 base/large
DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [48]. It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and utilizing Gradient-Disentangled Embedding Sharing. We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
BERT base/large
BERT is a well-established text embedding model using an encoder-only transformer trained with the masked language modelling objective [15]. We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased. While not state-of-the-art, BERT models remain widely used in embedding tasks.
BioClinicalBERT
CLIMBR-T-Base
CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million deidentified EHRs from Stanford Medicine with autoregressive next code prediction [13, 30]. It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base. The model has 141M parameters and allows for a context window of 496 codes. CLIMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [30]. It serves as a main baseline for our experiments to test specialized EHR models against general purpose text embeddings models for representing EHR records.
LLM-Embedding Model and CLIMBR-T-Base
To test whether the LLM-embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction. To this end we simply appended both embeddings. The resulting embeddings have dimensions 4.352 (GTE-Qwen2-7B) and 4.864 (LLM2Vec-Llama-3.1-8B).
Counts-based Model
Counts models have proven to be strong baselines for EHR prediction tasks [6, 12, 13]. The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept. We used the counts baseline introduced in [30] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
Based on the embeddings or the counts vectors given by the methods described above a classification head was trained and validated for each prediction task. For the embedding models we used a logistic regression head. For the counts-based model, we used a gradient boosting machine [39], which proved superior [30]. We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
4.5 Instructions for LLM-Embedding Models
The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts. Hence, we added simple instructions for each prediction task based on their respective instruction templates. For instance, for prediction of anemia we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”. The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons. However, to support task-specific instructions, we had to change this default behavior leading to 1.161.412 instead of 406.379 EHRs to encode resulting in longer processing times. The difference between 1.161.412 labels used in our experiments and the total number of labels of 1.178.665 (Table 2) is since some labels even have the same tasks and prediction time and are merged. We list all instructions in Table 8 and perform ablations testing the effect of the instructions.
4.6 Computational Setup and Running Times
All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems). Running the GTE-Qwen2-7B model with our serialization truncated at 4,096-tokens on an 8-GPU DGX system required approximately 20 hours. For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset. These issues were resolved by splitting the data into smaller batches, which introduced additional overhead. Additionally, we optimized the LLM2Vec code by removing an initial text pruning step that took approximately eight hours when using the full data. Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs.
4.7 Performance Results on 15 Prediction Tasks and Few-Shot Setting
Following the EHRSHOT benchmark, we evaluated all models across 15 prediction tasks under various few-shot settings. The benchmark includes a modular pipeline designed to execute key tasks, with the flexibility to optionally utilize a Slurm cluster for distributed execution. Running all steps within this pipeline ensures full reproducibility of results. Step four of the pipeline, which generates EHR representations with CLIMBR-T-Base and the counts-based model, was extended to incorporate our method for creating language model-based EHR representations. This adaptation allowed us to reuse significant portions of the existing code, including the task evaluation framework. Additionally, we implemented new functionality for EHR serialization and slightly modified other steps of the benchmark to accommodate our experimental setup. For instance, the label creation process was adjusted (step three) to enable task-specific instructions for the LLM-embedding models. All modifications have been documented and can be tracked in our public GitHub repository.
4.8 Effect of Different Contexts Sizes
We investigated the impact of varying context sizes in the LLM-embedding models. We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings. Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1.024, 2.048, and 8.192 tokens. Input tokens exceeding these thresholds were discarded. Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits. By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
4.9 Effect of Chunked Contexts
To further explore whether LLM-embedding models effectively interpret the entirety of their input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs. For this, the 4.096-token inputs were divided into smaller chunks of sizes 512, 1.024, and 2.048 tokens101010In our implementation we used these chunks sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models.. Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks. This approach aligns with the mean pooling applied to the last layer that we used to compute embeddings. In this setup, task-specific instructions, placed at the beginning of the inputs, were included only in the first chunk. This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
4.10 Ablations of EHR Serialization
To better understand the contribution of various components in the EHR serialization process to the performance of the LLM-embedding models, we conducted a series of ablation experiments. First, we assessed the impact of task-specific instructions by removing them entirely to determine their influence on the final embeddings. Subsequently, we performed additional ablations by systematically excluding specific components of the serialization. These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries. Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures. Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components. This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models leverage structured medical data for prediction tasks.
Appendix A Appendix
| Medical Concept | Semantic Codes | Unit | Min-Max Range | Normal Range | Formatting |
| Recent Body Metrics | |||||
| Body weight | LOINC/29463-7 | oz | 350-10000 | One decimal | |
| Body height | LOINC/8302-2 | inch | 5-100 | One decimal | |
| Body mass index / BMI | LOINC/39156-5 | kg/m2 | 10-100 | 18.5-24.9 | One decimal |
| Body surface area | LOINC/8277-6, SNOMED/301898006 | m2 | 0.1-10 | Two decimals | |
| Recent Vital Signs | |||||
| Heart rate | LOINC/8867-4, SNOMED/364075005, SNOMED/78564009 | bpm | 5-300 | 60-100 | Integer |
| Systolic blood pressure | LOINC/8480-6, SNOMED/271649006 | mmHg | 20-300 | 90-140 | Integer |
| Diastolic blood pressure | LOINC/8462-4, SNOMED/271650006 | mmHg | 20-300 | 60-90 | Integer |
| Body temperature | LOINC/8310-5 | °F | 80-120 | 95-100.4 | One decimal |
| Respiratory rate | LOINC/9279-1 | breaths/min | 1-100 | 12-18 | Integer |
| Oxygen saturation | LOINC/LP21258-6 | % | 1-100 | 95-100 | Integer |
| Recent Lab Results | |||||
| Hemoglobin | LOINC/718-7, SNOMED/271026005, SNOMED/441689006 | g/dL | 1-20 | 12-17 | One decimal |
| Hematocrit | LOINC/4544-3, LOINC/20570-8, LOINC/48703-3, SNOMED/28317006 | % | 10-100 | 36-51 | Integer |
| Erythrocytes | LOINC/789-8, LOINC/26453-1 | 106/uL | 1-10 | 4.2-5.9 | Two decimals |
| Leukocytes | LOINC/20584-9, LOINC/6690-2 | 103/uL | 1-100 | 4-10 | One decimal |
| Platelets | LOINC/777-3, SNOMED/61928009 | 103/uL | 10-1000 | 150-350 | Integer |
| Sodium | LOINC/2951-2, LOINC/2947-0, SNOMED/25197003 | mmol/L | 100-200 | 136-145 | Integer |
| Potassium | LOINC/2823-3, SNOMED/312468003, LOINC/6298-4, SNOMED/59573005 | mmol/L | 0.1-10 | 3.5-5.0 | One decimal |
| Chloride | LOINC/2075-0, SNOMED/104589004, LOINC/2069-3 | mmol/L | 50-200 | 98-106 | Integer |
| Carbon dioxide, total | LOINC/2028-9 | mmol/L | 10-100 | 23-28 | Integer |
| Calcium | LOINC/17861-6, SNOMED/271240001 | mg/dL | 1-20 | 9-10.5 | One decimal |
| Glucose | LOINC/2345-7, SNOMED/166900001, LOINC/2339-0, SNOMED/33747003, LOINC/14749-6 | mg/dL | 10-1000 | 70-100 | Integer |
| Urea nitrogen | LOINC/3094-0, SNOMED/105011006 | mg/dL | 1-200 | 8-20 | Integer |
| Creatinine | LOINC/2160-0, SNOMED/113075003 | mg/dL | 0.1-10 | 0.7-1.3 | One decimal |
| Anion gap | LOINC/33037-3, LOINC/41276-7, SNOMED/25469001 | mmol/L | -20-50 | 3-11 | Integer |
| Task | Prompt |
| Prefix | Given a patient’s EHR in Markdown format, retrieve relevant passages that answer the query: |
| Long Length of Stay | Will the patient stay in the hospital for more than 7 days |
| 30-day Readmission | Will the patient be readmitted to the hospital within 30 days |
| ICU Transfer | Will the patient be transferred to the intensive care unit |
| Thrombocytopenia | Has the patient thrombocytopenia |
| Hyperkalemia | Has the patient hyperkalemia |
| Hypoglycemia | Has the patient hypoglycemia |
| Hyponatremia | Has the patient hyponatremia |
| Anemia | Has the patient anemia |
| Hypertension | Has the patient hypertension |
| Hyperlipidemia | Has the patient hyperlipidemia |
| Pancreatic Cancer | Has the patient pancreatic cancer |
| Celiac | Has the patient celiac disease |
| Lupus | Has the patient lupus |
| Acute MI | Has the patient had an acute myocardial infarction |
| Chest X-Ray Findings | What are the chest x-ray findings of the patient |

### v3

@@
 Abstract
-Electronic Health Records (EHRs) offer rich potential for clinical prediction, yet their inherent complexity and heterogeneity pose significant challenges for traditional machine learning approaches.
-Domain-specific EHR foundation models trained on large collections of unlabeled EHR data have demonstrated promising improvements in predictive accuracy and generalization; however, their training is constrained by limited access to diverse, high-quality datasets and inconsistencies in coding standards and healthcare practices.
-In this study, we explore the possibility of using general-purpose Large Language Models (LLMs) based embedding methods as EHR encoders.
-By serializing patient records into structured Markdown text, transforming codes into human-readable descriptors, we leverage the extensive generalization capabilities of LLMs pretrained on vast public corpora, thereby bypassing the need for proprietary medical datasets.
-We systematically evaluate two state-of-the-art LLM-embedding models, GTE-Qwen2-7B-Instruct and LLM2Vec-Llama3.1-8B-Instruct, across 15 diverse clinical prediction tasks from the EHRSHOT benchmark, comparing their performance to an EHR-specific foundation model, CLIMBR-T-Base, and traditional machine learning baselines.
-Our results demonstrate that LLM-based embeddings frequently match or exceed the performance of specialized models, even in few-shot settings, and that their effectiveness scales with the size of the underlying LLM and the available context window.
-Overall, our findings demonstrate that repurposing LLMs for EHR encoding offers a scalable and effective approach for clinical prediction, capable of overcoming the limitations of traditional EHR modeling and facilitating more interoperable and generalizable healthcare applications.
-1 Introduction
-EHRs have become a widespread technology in modern healthcare, providing a comprehensive, longitudinal view of a patient’s health status [1].
-Machine learning methods can leverage this rich data to perform risk stratification and support clinical decision making [2, 3, 4].
+Electronic Health Records (EHRs) offer considerable potential for clinical prediction, but their complexity and heterogeneity present significant challenges for traditional machine learning methods.
+Recently, domain-specific EHR foundation models trained on large volumes of unlabeled EHR data have shown improved predictive accuracy and generalization.
+However, their development is constrained by limited access to diverse, high-quality datasets, and by inconsistencies in coding standards and clinical practices.
+In this study, we explore the use of general-purpose Large Language Models (LLMs) to encode EHR into high-dimensional representations for downstream clinical prediction tasks.
+We convert structured EHR data into markdown-formatted plain text documents by replacing medical codes with natural language descriptions.
+This enables the use of LLMs and their extensive semantic understanding and generalization capabilities as effective encoders of EHRs without requiring access to private medical training data.
+We show that LLM-based embeddings can often match or even surpass the performance of a specialized EHR foundation model, CLMBR-T-Base, across 15 diverse clinical tasks from the EHRSHOT benchmark.
+To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, a population distinct from that used to train CLMBR-T-Base.
+Notably, one of the tested LLM-based models achieves superior performance for disease onset, hospitalization, and mortality prediction, highlighting robustness to shifts in patient populations.
+Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and generalizable alternative to domain-specific models for clinical prediction.
+keywords:
+electronic health records, clinical prediction, machine learning, large language models, foundation models1 Introduction
+EHRs are now widely used in modern healthcare, providing comprehensive, longitudinal views of a patient’s health status [1].
+Machine learning methods can use this rich data for risk stratification and to support clinical decision-making [2, 3, 4].
 In recent years, researchers have explored a variety of prediction tasks based on EHRs, including hospital readmission [5, 6], length of hospital stay [6], sepsis onset detection [7, 8], mortality prediction [6, 9], discharge diagnoses [6], and heart failure outcomes [10].
-The overarching objective is to harness existing EHR data through machine learning to enhance clinical outcomes and reduce healthcare costs.
+The overarching goal is to harness EHR data using machine learning to improve clinical outcomes and reduce healthcare costs.
 However, machine learning on EHR data poses significant challenges due to its inherent complexity.
@@
 Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16].
-Treating EHR data as sequences of medical codes, enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13].
+Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13].
 However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data.
 Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
-In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural-language tasks [20].
+In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural language tasks [20, 21].
 This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications.
 Consequently, LLMs have demonstrated strong performance in extracting medical concepts [22], summarizing medical texts [23], and predicting medical outcomes [24] even in low-resource settings.
-However, most modern LLMs, such as GPT [24] or Llama [25], utilize a decoder-only transformer architecture, which complicates the generation of robust text representations.
+However, most modern LLMs, such as GPT [25] or Llama [26], use a decoder-only transformer architecture, which complicates the generation of robust text representations.
 To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [27, 28, 29, 30].
 Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
-In this study, we systematically evaluate whether general-purpose LLM-embedding models can effectively encode EHR records for 15 distinct clinical prediction tasks [30] (see Fig.
+In this study, we evaluate whether general-purpose LLM embedding models can effectively encode EHR data for clinical prediction [31] (see Fig.
 1).
-To this end, we first transform the EHR data into a concise text representation capturing the most relevant patient information at prediction time.
-We then use two LLM-embedding models, GTE-Qwen2-7B-Instruct [31, 29] and LLM2Vec-Llama3.1-8B-Instruct [26, 32], to generate EHR representations.
-The resulting embeddings serve as inputs to a logistic regression classifier, which is trained separately for each prediction task.
-We focus on evaluating the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness.
+To this end, we convert structured EHR records into plain text summaries that highlight key patient information at prediction time.
+Using two state-of-the-art LLM embedding models, GTE-Qwen2-7B-Instruct (GTE-Qwen2-7B) [32, 30] and LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B) [27, 33], we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark.
+We focus on the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness.
+Finally, we validate our approach on the UKB for predicting mortality, hospitalization and onset of 23 diseases [34] to assess its robustness and compare the generalizability with EHR-specific foundation models.
 2 Results
-2.1 EHR Database and Prediction Tasks
-We used the EHRSHOT database containing adult patients from the Stanford Health Care and Lucile Packard Children’s Hospital from 1990 to February 8th, 2023 [30].
-This dataset includes comprehensive EHR timelines for 6.739 patients, covering 921.499 visits, and 41.661.637 clinical events (Table 1).
-The database is part of a rigorous EHR benchmark containing 15 clinical prediction tasks grouped into four task groups.
-It also provides canonical dataset splits and provides publicly available code.
-Table 2 presents a detailed breakdown of the task groups, individual tasks, and the corresponding number of labels used in our experiments.
-Notably, a single patient may contribute multiple labels for a given prediction task, as relevant clinical events can recur over time.
-| Attribute | Train | Validation | Test | Total |
-| Num Events | 15,511,472 (37.2) | 13,005,205 (31.2) | 13,144,960 (31.6) | 41,661,637 (100.0) |
-| Num Visits | 339,504 (36.8) | 300,325 (32.6) | 281,670 (30.6) | 921,499 (100.0) |
-| Num Patients | 2,295 (34.1) | 2,232 (33.1) | 2,212 (32.8) | 6,739 (100.0) |
-| Num Female | 1,173 (34.1) | 1,142 (33.2) | 1,126 (32.7) | 3,441 (100.0) |
-| Num Male | 1,122 (34.0) | 1,090 (33.1) | 1,086 (32.9) | 3,298 (100.0) |
-| Age, mean SD | 59.6 17.8 | 58.8 18.0 | 59.5 18.0 | 59.3 17.9 |
-| American Indian | 14 (56.0) | 7 (28.0) | 4 (16.0) | 25 (100.0) |
-| Asian | 356 (34.1) | 347 (33.3) | 340 (32.6) | 1,043 (100.0) |
-| Black | 98 (32.9) | 105 (35.2) | 95 (31.9) | 298 (100.0) |
-| Pacific Islander | 23 (31.1) | 21 (28.4) | 30 (40.5) | 74 (100.0) |
-| Unknown | 518 (33.1) | 530 (33.9) | 515 (32.9) | 1,563 (100.0) |
-| White | 1,286 (34.4) | 1,222 (32.7) | 1,228 (32.9) | 3,736 (100.0) |
-| Hispanic | 374 (36.0) | 342 (32.9) | 322 (31.0) | 1,038 (100.0) |
-| Non-Hispanic | 1,921 (33.7) | 1,890 (33.2) | 1,890 (33.2) | 5,701 (100.0) |
+2.1 Experimental Setup
+Our primary analysis was conducted using the EHRSHOT database, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023.
+This dataset includes 921,499 visits and over 41.6 million clinical events [31], and serves as a standardized benchmark with 15 clinical prediction tasks organized into 4 task groups, along with predefined splits and publicly available code.
+Table 1 summarizes cohort statistics, and task details are shown in Table 2.
+Further information on task definitions and preprocessing is provided in Section 4.1.
+| Attribute | EHRSHOT | UK Biobank |
+| Num Patients | 6,739 | 387,464 |
+| Num Visits | 921,499 | 19,484,777 |
+| Num Events | 41,661,637 | 72,265,684 |
+| Num Female | 3,441 | 214,565 |
+| Num Male | 3,298 | 172,899 |
+| Age, mean SD | 59.3 17.9 | 56.78 8.11 |
+| American Indian | 25 | 0 |
+| Asian | 1,043 | 8,659 |
+| Black | 298 | 5,751 |
+| Pacific Islander | 74 | 0 |
+| Unknown | 1,563 | 7,202 |
+| White | 3,736 | 365,888 |
+| Hispanic | 1,038 | - |
+| Non-Hispanic | 5,701 | - |
 | Attribute | Train Labels (Positive) | Valid Labels (Positive) | Test Labels (Positive) | Total Labels (Positive) |
-| Operation Outcomes | ||||
+| Operational Outcomes | ||||
 | Long Length of Stay | 2,569 (681) | 2,231 (534) | 2,195 (552) | 6,995 (1,767) |
@@
 | Chest X-Ray Findings | 7,481 (4,771) | 9,366 (6,032) | 9,428 (6,400) | 26,275 (17,203) |
-2.2 EHR Text Serialization
-To utilize an LLM-embedding model for EHR representation, the patient records were encoded into a structured text based on the widely used Markdown format.
-Due to runtime constraints, the maximum serialization length was limited to 4.096 tokens (approximately 16.000 characters).
-This constraint informed the serialization strategy, prioritizing the inclusion of recent data to ensure critical medical information was preserved.
-The serialized record is divided into clearly labeled sections (see example in Fig.
-2).
-All dates were normalized relative to a reference prediction date of January 1st, 2024, explicitly stated at the beginning.
-This is followed by the patient’s basic demographic information.
-Approximately 65% of the recorded values were time-series data of Logical Observation Identifiers Names and Codes (LOINC) concepts.
-To reduce the volume of this data, the serialization focuses on 24 frequently recorded concepts, grouped into three primary categories: Body Metrics, Vital Signs, and Lab Results.
-For each selected concept, the last three recorded values were included, along with the corresponding units and a classification as low, normal, or high, where applicable.
-The serialization then summarizes all patient visits, including visit type, time and duration, followed by any event not associated with specific visits.
-Detailed visit records are then provided in descending chronological order, beginning with the most recent visit.
-Each visit entry is further categorized into conditions, medications, and procedures for improved clarity and utility.
-As a result, the text serialization ensures a structured and concise representation of EHR data, facilitating efficient processing by LLM-based models.
-2.3 LLM-Embedding Models and Baselines
-We focus on two LLM-embedding models as part of our experiments to explore their effectiveness in representing EHR data.
-The first model, GTE-Qwen2-7B-instruct (GTE-Qwen2-7B), is based on the Qwen2-7B LLM [31] and incorporates bidirectional attention and contrastive learning to enhance embedding tasks [29].
-The second model, LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B), is built on the Llama-3.1-8B Instruct architecture [32] and employs similar optimization techniques to improve embeddings [26].
-Both models were trained using instructions for the embeddings task.
-Hence, we added a simple prompt for each task, e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia?” (see Table 8).
-Our primary goal was to assess how these LLM-based models, trained on publicly available text data, perform in representing EHRs compared to an EHR-specific foundation model.
-For this, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 30].
-For each model, we used a logistic regression classification head trained on the training split, with hyperparameters tuned on the validation split.
-Additionally, we included a counts-based baseline using a Gradient Boosted Machine (GBM) model, which has demonstrated superior performance over logistic regression for EHR tasks [30].
+For external validation, we used the UKB, a population-based cohort comprising 502,489 UK participants [35, 36].
+This allowed us to evaluate the generalizability of our approach across healthcare systems, especially since the CLMBR-T-Base model was trained on data from the same hospital system as EHRSHOT.
+Due to structural differences between the datasets, not all EHRSHOT tasks were transferable to the UKB (see Section 4.8).
+We focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 chronic diseases [34].
+The processed UKB subset used in our study consisted of 387,464 patients, approximately 19.5 million visits, and over 72 million clinical events.
+Table 6 provides task descriptions and label distributions.
+To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 4,096 tokens, prioritizing more recent data.
+For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key LOINC concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high).
+These were followed by a summary of patient visits and reverse-chronological listings of visit-level events such as conditions, medications, and procedures.
+An example EHR text serialization is shown in Fig.
+2, and a detailed explanation appears in Section 4.2.
+For the UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations.
+We evaluated two instruction-tuned LLM embedding models, GTE-Qwen2-7B [32, 30] and LLM2Vec-Llama-3.1-8B [33, 27], using task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 8).
+Our primary objective was to assess how general-purpose LLM-based embedding models, trained on public text data, perform in encoding EHRs for clinical prediction tasks.
+For comparison, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 31].
+For each model, we computed patient-level embeddings and trained a logistic regression classifier on the training split, with hyperparameters selected using a validation set.
+As a baseline, we also included a Gradient Boosted Machine (GBM) classifier trained on count-based features, which has previously demonstrated strong performance on EHR tasks [31].
+See Section 4.4 for additional details on model architectures.
 | Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
@@
 | ClinicalBERT | |||||
-2.4 Performance Results on 15 Prediction Tasks
-The performance results measured by the Area Under the Receiver Operating Characteristic Curve (AUROC) for all training and validation examples are presented in Table 3.
-The GTE-Qwen2-7B model demonstrated superior performance over CLIMBR-T-Base in three of the four task categories: operation outcomes, lab test result prediction, and assignment of new diagnoses.
-However, in predicting chest X-ray findings, the EHR foundation model (CLIMBR-T-Base) outperformed GTE-Qwen2-7B.
-The LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance compared to GTE-Qwen2-7B, surpassing CLIMBR-T-Base only in the assignment of new diagnoses task.
-Both LLM-based embedding models clearly outperformed the counts-based baseline with a gradient-boosted machine head, which is traditionally regarded as a strong baseline for EHR tasks [12, 13].
-The only exception was in the assignment of new diagnoses, where the counts baseline marginally outperformed GTE-Qwen2-7B.
-We also combined the CLMBR-T-Base embedding with the LLM representations by simple concatenation to test whether they encode orthogonal information.
-The combined embeddings led to a considerable performance boost for both models with an average AUROC performance of 0.801 (0.777 - 0.826) instead of 0.774 (0.749 - 0.798) for GTE-Qwen2-7B and 0.779 (0.753 - 0.805) instead of 0.742 (0.714 - 0.769) for LLM2Vec-Llama-3.1-8B.
-The smaller GTE-Qwen2-1.5B model only performed slightly worse than GTE-Qwen2-7B with an average performance of 0.757 (0.732 - 0.783) compared to 0.774 (0.749 - 0.798) making it a performant model choice for EHR embeddings.
-LLM2Vec-Llama-2-1.3B on the other hand performed significantly worse than the larger model which could be due to the older Llama 2 architecture.
-Fig.
-3 illustrates the relationship between model size and average performance across all tasks, suggesting the trend that larger models tend to perform better.
-Among the evaluated models, CLIMBR-T-Base proved to be the most efficient in terms of performance relative to model size.
-It is noteworthy, however, that the LLM-based embedding models were not specifically optimized for this EHR data modality, which may partially explain their performance gap.
-2.5 Performance Results in Few-Shot Setting
+| Model | Mortality prediction | Operational Outcomes (Hospitalization) | Assignment of New Diagnoses | Macro Avg. Across Task Groups |
+| LLM-Embedding Models | ||||
+| GTE-Qwen2-7B | ||||
+| LLM2Vec-Llama-3.1-8B | ||||
+| Baselines [31] | ||||
+| CLMBR-T-Base | ||||
+| Counts-based + GBM | ||||
+2.2 General-Purpose LLM Embeddings Rival Domain-Specific EHR Models
+Using all available training and validation examples, the GTE-Qwen2-7B model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3).
+In contrast, CLMBR-T-Base achieved higher performance on tasks related to operational outcomes and chest X-ray findings.
+LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance than GTE-Qwen2-7B, exceeding CLMBR-T-Base only in the assignment of new diagnoses task group.
+Both LLM-based models were often on par with or outperformed the GBM-based baseline, which is considered a strong baseline for EHR prediction tasks [12, 13].
+To assess whether the models encode orthogonal information, we concatenated CLMBR-T-Base embeddings with LLM-based embeddings and observed substantial performance improvements for both LLM models.
+The combined embeddings achieved an average area under the receiver operating characteristic curve (AUROC) of 0.786 (0.762–0.811) for GTE-Qwen2-7B, up from 0.755 (0.730–0.780), and 0.773 (0.747–0.799) for LLM2Vec-Llama-3.1-8B, up from 0.734 (0.707–0.762).
+The smaller GTE-Qwen2-1.5B model achieved an average performance of 0.748 (0.722–0.774), only marginally lower than its larger counterpart, suggesting it is a viable lightweight alternative for EHR embeddings.
+In contrast, LLM2Vec-Llama-2-1.3B performed notably worse than the larger version, potentially due to limitations of the older Llama 2 architecture.
+Figure 3 shows a general trend that larger models tend to yield better performance across tasks.
+Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance.
+Nonetheless, it is important to note that the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
+On the UKB dataset, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all three prediction task groups (see Table 4).
+LLM2Vec-Llama-3.1-8B surpassed CLMBR-T-Base only in the task group related to the assignment of new diagnoses.
+The performance drop of CLMBR-T-Base was anticipated, as the UKB represents an out-of-domain dataset with numerous disease codes not included in the model’s original vocabulary.
+Although we applied a careful mapping of UKB Observational Medical Outcomes Partnership (OMOP) concepts to those supported by CLMBR-T-Base to improve compatibility and predictive performance, many concepts remained unmapped due to CLMBR-T-Base’s limited vocabulary.
+This limitation underscores a key challenge in adapting domain-specific models to external datasets, even with considerable manual effort.
+In contrast, the LLM-based models required only a mapping from clinical concepts to their natural language descriptions, which were directly incorporated into the Markdown-based serialization.
+This simplified integration process substantially reduced implementation overhead and emphasized the strong generalization ability of LLM-based encodings across diverse coding systems and healthcare domains.
+2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
 To evaluate model performance with limited training data, we conducted experiments in a few-shot setting using small numbers of training examples.
@@
 Fig.
-4 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples (see Fig.
-6 for AUPRC results).
-Both LLM-embedding models demonstrated strong results in the few-shot setting, indicating that their pretraining on general text can be successfully transferred to serialized EHR data.
-Notably, GTE-Qwen2-7B consistently outperformed CLIMBR-T-Base across all training example sizes for predicting lab test results and assigning new diagnoses.
-For operational outcomes, a minimum of 32 training examples was required for GTE-Qwen2-7B to surpass the performance of the EHR-specific foundation model.
-In contrast, LLM2Vec-Llama-3.1-8B showed weaker performance in the few-shot setting.
-It only outperformed CLIMBR-T-Base in a limited number of scenarios, primarily for the assignment of new diagnoses.
-Both LLM-embedding models consistently achieved improvements over the counts-based baseline across most configurations, underscoring the impact of their pretraining on general text.
-Consistent with previous findings [30], the largest performance gains over the counts-based baseline were observed at intermediate training example sizes.
-As the number of training examples increased, the advantage of pretrained LLM-based models decreased, highlighting the diminishing returns of foundation models as more labeled data becomes available.
-2.6 Effect of Different Contexts Sizes
-We evaluated the impact of varying context sizes on the performance of the LLM-embedding models by testing GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B with input lengths of 512, 1.024, 2.048, 4.096 (the default), and 8.192 tokens.
-The objective was to determine whether longer input sequences enhance performance and to assess if the models effectively leverage additional historical data.
-As summarized in Table 4 and Fig.
-5, the two models exhibited distinct behaviors.
+5 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples.
+LLM embedding models performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data.
+Notably, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all training sizes for lab test result prediction and assignment of new diagnoses.
+The strong results for lab test predictions indicate that the inclusion of the last three lab values in the EHR serialization provides highly predictive features, which GTE-Qwen2-7B effectively encodes.
+In contrast, LLM2Vec-Llama-3.1-8B showed comparatively weaker performance in the few-shot regime.
+It only surpassed CLMBR-T-Base in a limited subset of tasks, primarily those involving the assignment of new diagnoses.
+Nonetheless, both LLM-based models consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining.
+Consistent with previous findings [31], the largest performance improvements over the baseline were observed at intermediate training set sizes.
+As the number of labeled data samples increased, the relative advantage of pre-trained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes.
+The area under the precision-recall curve (AUPRC) results are shown in Fig.
+7, with task-specific AUROC and AUPRC results provided in Fig.
+8 and Fig.
+9 in the Appendix.
+In contrast to EHRSHOT, the LLM-based methods outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset.
+For the assignment of new diagnoses, which includes 23 tasks, both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed CLMBR-T-Base across all sample sizes (see Fig.
+10 in the Appendix).
+For hospitalization and mortality prediction, CLMBR-T-Base showed stronger performance in the very low-data regime.
+However, with more samples the performance of the LLM embedding models increased and GTE-Qwen2-7B outperformed CLMBR-T-Base.
+Figure 5 presents the mean AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences.
+On EHRSHOT, GTE-Qwen2-7B generally matched CLMBR-T-Base in performance, while LLM2Vec-Llama-3.1-8B fell between these models and the count-based baseline.
+On the UKB dataset, LLM-based models and CLMBR-T-Base performed comparably with very limited training data of up to four examples.
+With additional training data, GTE-Qwen2-7B surpassed both CLMBR-T-Base and LLM2Vec-Llama-3.1-8B, while the latter closely approached the performance of CLMBR-T-Base.
+These findings underscore the capability of general-purpose LLM embeddings to rival or exceed the performance of domain-specific models, particularly in settings characterized by domain shift or limited labeled data.
+Detailed task-specific results are available in Fig.
+12 and Fig.
+13 in the Appendix.
 | Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
 | GTE-Qwen2-7B | |||||
-| 8.192 context size | |||||
-| 2.048 context size | |||||
-| 1.024 context size | |||||
-| 512 context size | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| 8.192 context size | |||||
-| 2.048 context size | |||||
-| 1.024 context size | |||||
-| 512 context size |
-For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4.096 tokens, with only a slight decline observed when using 2.048 or 1.024 tokens.
-This suggests that the most relevant information is contained within the first 1.024 tokens and that the model can accurately extract these details even when presented within a longer context.
-In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2.048 tokens and a moderate gain at 1.024 tokens, indicating that the model struggles to effectively handle longer sequences up to 4.096 tokens.
-Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [33].
-Furthermore, extending the context to 8.192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs.
-This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
-2.7 Effect of Chunked Contexts
-To investigate whether the models can process the full 4.096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1.024, and 2.048 tokens.
-For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 5).
-For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4.096-token input remains effectively used even when segmented.
-Notably, using 512-token chunks yielded an overall AUROC performance of 0.735, compared to 0.672 when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints.
-In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes.
-This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2.048 tokens.
-| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
-| 2 x 2.048 tokens chunks | |||||
-| 4 x 1.024 tokens chunks | |||||
-| 8 x 512 tokens chunks | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| 2 x 2.048 tokens chunks | |||||
-| 4 x 1.024 tokens chunks | |||||
-| 8 x 512 tokens chunks |
-| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
+| generic instructions | |||||
 | no instructions | |||||
@@
 | LLM2Vec-Llama-3.1-8B | |||||
+| generic instructions | |||||
 | no instructions | |||||
@@
 | no procedures |
-2.8 Ablations of EHR Serialization
-To assess the impact of different components of the EHR serialization, we conducted a series of ablation studies for both LLM-embedding models, as summarized in Table 6.
-The removal of instructions had a notable effect on performance for both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, leading to a decrease from 0.774 (0.749 - 0.798) to 0.725 (0.700 - 0.750) and from 0.742 (0.714 - 0.769) to 0.727 (0.700 - 0.755) in averaged AUROC, respectively.
-The most significant drop was observed in tasks related to operational outcomes and lab result prediction, suggesting that the instructions played a critical role in guiding the model to focus on relevant clinical information.
-The inclusion of aggregated information, comprising four body metrics, six vital signs, and 14 lab values, also had a major impact on predictive performance.
-Removing this information substantially reduced the accuracy of lab test predictions, with performance dropping from 0.867 (0.860 - 0.874) to 0.713 (0.701 - 0.726) for GTE-Qwen2-7B and from 0.777 (0.766 - 0.787) to 0.705 (0.692 - 0.718) for LLM2Vec-Llama-3.1-8B.
-This highlights the importance of the most recent lab values in predicting future lab results.
-The effect was particularly pronounced for GTE-Qwen2-7B in predicting conditions such as thrombocytopenia, hyperkalemia, and hyponatremia.
-However, for hypoglycemia and anemia, GTE-Qwen2-7B performed on par with CLIMBR-T-Base, even though the lab values were still part of the serialized representation (see Fig.
-7 and Fig.
-8 for task specific performance).
-Interestingly, the removal of aggregated information led to slight performance improvements in predicting operational outcomes and the assignment of new diagnoses, suggesting that in some cases, a more focused representation may be beneficial.
-Visit-related information proved to be particularly critical for operational outcome prediction, as removing it led to a performance drop from 0.844 (0.821 - 0.867) to 0.752 (0.728 - 0.776) for GTE-Qwen2-7B and from 0.787 (0.762 - 0.812) to 0.734 (0.709 - 0.759) for LLM2Vec-Llama-3.1-8B.
-Notably, the removal of visit information resulted in a substantial performance increase for LLM2Vec-Llama-3.1-8B predicting lab values, from 0.777 (0.766 - 0.787) to 0.862 (0.855 - 0.869).
-This suggests that the model struggles to extract relevant information when presented with a large volume of contextual data, a challenge that GTE-Qwen2-7B handles more effectively.
-On the other hand, the removal of demographic information, conditions, medications, and procedures had only a minor effect on performance, indicating that these features were either redundant or less relevant for the prediction tasks.
-While the influence of aggregated information and visit details was clear for predicting operational outcomes and lab results, the ablation studies showed less impact on tasks related to new diagnoses and chest X-ray prediction.
-This suggests that these latter task groups rely more on a broader set of contextual factors rather than any single component of the serialization.
+2.4 Effect of Serialization Components on LLM-Based EHR Embedding Performance
+To evaluate the contribution of individual serialization components, we conducted ablation studies for both LLM embedding models (see Table 5).
+Replacing task-specific instructions with a generic prompt led to a marked performance decline, reducing average AUROC from 0.755 (0.730-0.780) to 0.739 (0.713-0.765) for GTE-Qwen2-7B and from 0.734 (0.707-0.762) to 0.724 (0.696-0.752) for LLM2Vec-Llama-3.1-8B.
+Removing the instruction entirely caused a further drop for GTE-Qwen2-7B to 0.724 (0.698-0.749), whereas LLM2Vec-Llama-3.1-8B showed minimal change, suggesting greater reliance on instruction prompting by GTE-Qwen2-7B.
+The most significant drop was observed in lab result prediction tasks, highlighting the importance of guidance via task-specific instructions.
+Aggregated measurements (body metrics, vital signs, and lab values) were crucial, especially for lab test prediction, where their removal reduced AUROC from 0.865 (0.858-0.873) to 0.714 (0.701-0.726) for GTE-Qwen2-7B and from 0.778 (0.768-0.789) to 0.706 (0.693-0.719) for LLM2Vec-Llama-3.1-8B.
+This underscores the predictive value of recent lab results and their effective encoding by the models.
+For conditions like thrombocytopenia, hyperkalemia, and hyponatremia, the performance drop was especially pronounced without this information for GTE-Qwen2-7B (see Fig.
+8).
+Interestingly, removing aggregated values slightly improved new diagnosis prediction for GTE-Qwen2-7B, suggesting a more focused input may benefit some tasks.
+Visit-level information had a strong influence on operational outcome tasks.
+Its removal led to AUROC reductions from 0.771 (0.747-0.795) to 0.752 (0.728-0.776) for GTE-Qwen2-7B and from 0.753 (0.727-0.778) to 0.734 (0.709-0.759) for LLM2Vec-Llama-3.1-8B.
+However, for LLM2Vec-Llama-3.1-8B, removing visit data unexpectedly improved lab result prediction, suggesting possible difficulty in filtering relevant signals from verbose input.
+Removing demographics, conditions, medications, and procedures had only minor effects, indicating these features were either redundant or contributed little predictive signal in isolation.
+While aggregated information and visit data were key drivers for lab and operational outcomes, prediction of new diagnoses and chest X-ray findings was less sensitive to individual serialization components.
+This suggests these tasks rely on broader contextual patterns rather than specific structured elements.
+2.5 Effect of Temporal Scope and Context Length on Embedding Effectiveness
+To assess model sensitivity to input length and recency, we performed ablations on temporal window, context size, and chunking using EHRSHOT data.
+Both LLM embedding models showed improved performance when restricted to more recent data, with the 1-month window outperforming longer histories.
+This suggests that recent clinical information may carry the most predictive signal, while older records may contribute less relevant or potentially distracting details (see Fig.
+6 and Section A.2).
+The models differed in optimal context size: GTE-Qwen2-7B achieved the best performance at the maximum 4,096-token context, whereas LLM2Vec-Llama-3.1-8B performed best at 2,048 tokens (Fig.
+6 and Table 10).
+Both models suffered large drops at 512 tokens, a common constraint in older language models [37].
+To test whether models processed the entire input in context, we evaluated chunked context representations, where fixed-size segments (e.g., 512–2,048 tokens) of a 4,096-token input were embedded separately and averaged (see Table 11).
+GTE-Qwen2-7B showed little degradation from chunking, indicating that simple averaging can be sufficient to preserve performance even under token limits.
+In contrast, LLM2Vec-Llama-3.1-8B benefited from chunking, suggesting it struggles with longer contiguous sequences and prefers shorter inputs.
+These results highlight differences in model-specific sequence processing and suggest practical approaches for adapting to context length constraints.
+Full tables and additional details are available in Section A.3 and Section A.4.
 3 Discussion
-Our study demonstrates that general-purpose LLM-embedding models, originally pre-trained on extensive natural language corpora, can be repurposed as robust foundation models for EHR prediction tasks.
-Specifically, models like GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B not only surpassed a strong counts-based baseline but, in several clinical domains, matched or even exceeded the performance of a dedicated EHR foundation model (CLIMBR-T-Base) [13, 30].
-Notably, this is despite CLIMBR-T-Base being trained on data from the same hospital system as the EHRSHOT database, underscoring the strong generalization capabilities of LLM-based embeddings.
-This is especially significant given the inherent challenges of limited and heterogeneous EHR data.
-Across 15 diverse clinical tasks, including few-shot scenarios, our findings indicate that the transferable knowledge acquired during large-scale text pretraining enables these models to effectively capture and encode complex clinical information.
-Furthermore, our analysis shows that performance scales with both model size and architectural improvements, suggesting that future advancements in general-purpose LLMs may further enhance their applicability in healthcare settings.
-In contrast, traditional EHR-specific models face scaling limitations due to restricted data availability and the necessity for specialized architectures [34].
-Collectively, these results highlight that repurposing LLMs as embedding generators offers a powerful, flexible, and scalable alternative for clinical prediction tasks.
-Our experiments suggest that several key factors contribute to the success of LLM-embedding models for encoding EHR data.
-First, providing clear, task-specific instructions within the serialized EHR text was crucial, as it guided the models to focus on the most clinically relevant sections, particularly benefiting tasks that require attention to specific input segments.
-This design leverages the general-purpose pretraining and further instruction-based fine-tuning of LLM-embedding models [26, 29], enabling them to identify and extract meaningful patterns from heterogeneous and complex EHR inputs.
-Second, the inclusion of aggregated clinical data, especially semantic LOINC-coded events and detailed visit-related information, was strongly associated with improved performance on operational outcomes and lab test predictions.
-In contrast, other components of the EHR did not show as clear a relationship with predictive accuracy.
-The effective handling of context size also plays a significant role.
-While models like GTE-Qwen2-7B can fully utilize contexts up to 4,096 tokens, our findings show that performance degrades with overly short or excessively long inputs.
-Short contexts risk omitting recent, crucial clinical encounters, whereas overly long contexts can dilute important token-level details.
-In cases where models face challenges with longer inputs such as LLM2Vec-Llama-3.1-8B our chunking experiments indicate that segmenting the input and averaging the resulting embeddings can mitigate this issue.
-Finally, combining LLM embeddings with those from an EHR-specific foundation model (CLMBR-T-Base) led to a considerable boost in performance.
-This improvement suggests that LLMs capture orthogonal and complementary clinical information.
-One explanation for this effect could be the limited context size of CLMBR, restricted to 496 medical codes, which prevents it from fully leveraging all information present in our EHR serialization [13].
-Alternatively, the boost may stem from the general-purpose pretraining on vast text corpora, which provides LLMs with domain-agnostic knowledge that remains untapped in models trained exclusively on EHR data.
-Unlike count-based models and specialized EHR foundation models, LLM-based models are agnostic towards specific coding systems and data formats.
-Traditional EHR models depend on predefined vocabularies such as SNOMED, LOINC, and ICD, which restricts their applicability across diverse healthcare settings.
-In contrast, LLMs operate on raw textual representations, enabling them to interpret any code that can be mapped to a human-readable form.
-This flexibility is particularly valuable given the persistent challenges in EHR data interoperability, stemming from variable coding practices, privacy constraints, and regulatory limitations, which often hinder the aggregation of large, standardized datasets for model pretraining [35].
+Our study shows that general-purpose LLM embedding models, pre-trained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data.
+Models such as GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 31].
+This is notable, as CLMBR-T-Base was trained on data from the same health system as EHRSHOT.
+The strong performance of the LLM-based encoders on the UKB dataset highlights their generalization capabilities.
+GTE-Qwen2-7B consistently outperformed the domain-specific foundation model with at least 16 training examples, while LLM2Vec-Llama-3.1-8B achieved similar results.
+These results are particularly notable given the challenges posed by EHR data, including limited size, heterogeneity, and domain shift.
+Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pre-training enables LLMs to extract clinically meaningful patterns from serialized EHRs.
+Model performance improved with larger model sizes and more advanced architectures, suggesting further gains are likely with future LLM development.
+In contrast, domain-specific EHR models remain constrained by the scarcity and fragmentation of large, diverse clinical datasets.
+Overall, our findings support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
+Our experiments suggest that several key factors contribute to the success of LLM embedding models for encoding EHR data.
+First, providing clear, task-specific instructions in the serialized EHR text was crucial, guiding models to focus on clinically relevant sections and improving tasks requiring targeted attention.
+This takes advantage of the instruction-tuned nature of LLM embedding models [27, 30], enabling them to extract meaningful patterns from heterogeneous EHR inputs.
+Second, aggregating structured data, particularly recent lab results and visit-level information, substantially improved performance, especially for lab test and operational outcome prediction.
+In contrast, components like demographics, medications, and condition codes contributed less consistently to model accuracy.
+Context size was also critical: GTE-Qwen2-7B effectively handled up to 4,096 tokens, while both models performed poorly with very short or overly long inputs.
+Intermediate time windows, especially those covering one month prior to prediction, yielded optimal performance, suggesting that recent clinical data carries more predictive value than long-term history for many tasks.
+For instance, lab values reflect dynamic physiological states and may lose relevance outside of a narrow temporal window.
+This indicates that LLM embeddings are particularly effective at encoding temporally focused, task-relevant clinical signals.
+Notably, GTE-Qwen2-7B demonstrated greater robustness to extended inputs, highlighting its stronger capacity to extract salient information from longer contexts.
+Finally, combining LLM embeddings with CLMBR-T-Base representations improved overall performance, suggesting that these models capture complementary dimensions of patient history.
+This synergy may stem from CLMBR-T-Base’s limited code-based input window of 496 medical codes [13], or from the broader world knowledge encoded in LLMs that is absent in domain-specific models.
+Unlike count-based models and specialized EHR foundation models, LLM-based methods are agnostic to coding systems and data formats.
+Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit their applicability across healthcare settings.
+For instance, CLMBR-T-Base processes only 65,536 codes (26,249 unique), selected based on their entropy from the EHRSHOT dataset [31].
+This restriction may lead to the exclusion of clinically important information when applied to external datasets with unseen codes.
+In contrast, LLMs operate on textual inputs, allowing them to interpret any clinical code mapped to a human-readable description.
+This flexibility is particularly valuable given the ongoing challenges in EHR interoperability due to varying coding practices, privacy constraints, and regulatory barriers, which often hinder the aggregation of large, standardized datasets for model pre-training [38].
 Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [34].
-This capacity mitigates the limitations of count-based models that typically filter out infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [37].
-Additionally, the ability of LLMs to process both structured and unstructured data paves the way for integrated, multimodal representations that could further enhance clinical decision support systems [38].
-Our findings highlight that LLM-embedding methods merge the strengths of count-based models, specialized EHR models, and the expansive generalization capabilities of large language models.
-While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are inherently limited by their reliance on predefined vocabularies and constrained training data.
-In contrast, LLMs can process any textual representation, enabling them to effectively handle rare or unseen codes, although they typically produce unstructured outputs that challenge clinical grounding and calibration [23].
-LLM-embedding methods bridge this gap by leveraging the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks.
-By transforming EHR data into structured text and embedding it using general-purpose LLMs, we enable the application of logistic regression or other conventional machine learning classifiers to make clinically meaningful predictions.
-This approach preserves the calibration advantages of traditional models while benefiting from the rich contextual understanding of LLMs.
-3.1 Limitations
+This mitigates the limitations of count-based models that often exclude infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [39].
+Finally, the ability to represent both structured and unstructured information makes LLMs well-suited for future multimodal EHR applications, including those that integrate clinical notes, diagnostic reports, and imaging metadata [40].
+Our findings highlight that LLM embedding methods combine the strengths of count-based models, specialized EHR models, and the generalization capabilities of LLMs.
+While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are limited by predefined vocabularies and constrained training data.
+In contrast, LLMs can process any textual representation, enabling them to handle rare or unseen codes, though they often produce unstructured outputs that complicate clinical grounding and calibration [24].
+LLM embedding methods bridge this gap by using the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks.
+By converting EHR data into structured text and embedding it with general-purpose LLMs, we enable the use of standard classifiers such as logistic regression to generate clinically meaningful predictions.
+This approach retains the calibration strengths of traditional models while benefiting from the rich contextual understanding of LLMs.
 This study has several limitations.
-First, our approach relies on a subjectively designed EHR serialization that we deemed to capture the most medically relevant information.
-This design choice may introduce bias when comparing LLM-based embedding models with dedicated EHR foundation models that operate on raw data.
-In addition, the performance of the LLM-embedding models is sensitive to the specific content and instructions provided in the serialized text, which may limit reproducibility and generalization ability.
-Although these models achieve competitive predictive performance across diverse clinical tasks, including few-shot scenarios, their substantially larger parameter counts lead to longer computation times and higher resource usage.
-Moreover, by relying on LLM embedding methods, we must train a downstream classifier from scratch, thereby forgoing the inherent zero-shot or few-shot prompting capabilities of LLMs.
-Finally, our serialization was limited to 4,096 tokens to manage runtime constraints, potentially omitting valuable long-range historical information, and our evaluation on a single institutional dataset may limit broader applicability across diverse healthcare systems.
-3.2 Future work
-Future research should explore serialization‐free approaches that allow LLMs to process raw EHR data directly, thereby reducing potential biases introduced by manual text transformation.
+First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information, which may introduce bias when comparing against models trained on raw data.
+Additionally, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization.
+While the models achieve strong predictive accuracy across tasks, including few-shot settings, their large parameter counts lead to longer computation time and higher resource demands.
+Moreover, the use of LLM embeddings requires training a separate downstream classifier, foregoing the native zero-shot or few-shot capabilities of LLMs.
+We also limited EHR serialization to 4,096 tokens due to runtime constraints, which may exclude relevant long-term historical data.
+Finally, evaluation was conducted on only two datasets, which may limit the generalizability of our findings across diverse healthcare systems.
+Future research should explore serialization-free approaches that allow LLMs to process raw EHR data directly, reducing biases introduced by manual text transformation.
 Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training.
-It will also be important to develop strategies to extend the effective context window beyond 4.096 tokens to capture more comprehensive patient histories.
-Moreover, investigating techniques for distilling large LLMs into smaller, more efficient models may enhance their practical applicability in clinical settings.
-Finally, expanding evaluations to multi-institutional datasets and examining how complementary insights from both domain-specific EHR models and general-purpose LLMs can be synergistically combined will be critical for advancing the development of robust, scalable EHR foundation models.
+Extending the effective context window beyond 4,096 tokens will be important for capturing comprehensive patient histories.
+Additionally, developing techniques to distill large LLMs into smaller, more efficient models could improve their practical use in clinical settings.
+Expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
 4 Methods
-4.1 EHR Database and Prediction Task
-The EHR data utilized in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [30].
+4.1 EHRSHOT Database and Prediction Task
+The EHR data used in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [31].
 We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement.
-This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8th, 2023.
+This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8, 2023.
 Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table.
-We used the official ehrshot-benchmark repository111Github repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments enabling to us to build on existing functionalities and to facilitate comparisons with prior methods.
+We used the official ehrshot-benchmark repository111GitHub repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments, enabling us to build on existing functionalities and to facilitate comparisons with prior methods.
 The benchmark uses the Framework for Electronic Medical Records (FEMR) 222GitHub repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data.
 All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark.
-The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical predictions tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [30].
-Task labels are derived from clinical events, resulting in significant variations in task-specific sample sizes.
+The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [31].
+Task labels are derived from clinical events; hence, a single patient could contribute multiple labels per task, resulting in significant variations in task-specific sample sizes.
 For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses.
@@
 Testing is always performed on the full set of examples.
-The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and gradient boosting machines [39].
-Performance is reported using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC).
+The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and GBMs [41].
+Performance is reported using the AUROC and the AUPRC.
 For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [31].
@@
 4.2 EHR Text Serialization
-To leverage LLM-embedding models for representing EHR records, we serialized the records into textual formats.
-We had to limit the length of the serializations to 4.096 tokens (approximately 16.000 characters) to carry out all experiments on the available computing infrastructure (see Section 4.6).
-Two experimental runs were conducted with serializations extending up to 8.192 tokens.
+To use LLM embedding models for representing EHR records, we serialized the records into textual formats.
+We limited the serialization length to 4,096 tokens (approximately 16,000 characters) due to computational constraints (see Section 4.6).
+Two experimental runs were conducted with serializations extending up to 8,192 tokens.
 The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text.
 This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted.
-To convert the visits and clinical events in the EHRSHOT dataset into text, we leveraged the semantic information embedded in the dataset.
+To convert the visits and clinical events in the EHRSHOT dataset into text, we used the semantic information embedded in the dataset.
 Each clinical event was labeled using the format “ontology/code”.
 EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated.
-An analysis was performed to evaluate the utilization of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2.968 labels.
+An analysis was performed to evaluate the use of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2,968 labels.
 We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type.
 We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type).
-Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontology.
-Cancer Modifier code contained UICC cancer stages that we parsed manually.
+Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontologies.
+Cancer Modifier codes contained UICC cancer stages that we parsed manually.
 For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx.
-We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful description.
+We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful descriptions.
 Various approaches exist for serializing structured data, including row-by-row serialization [42], template-based methods [43], or structured data formats like JSON, HTML, and Markdown [44].
@@
 To avoid duplicates, we merged synonymous codes (see Table 7).
-The last three values of each concept are given, and we filtered implausible values.
+The last three values of each concept are presented, and we filtered implausible values.
 To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 7).
@@
 To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records.
-Also, due to computational constraints, we evaluated only three variants of our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig.
+Also, due to computational constraints, we evaluated only three alternatives to our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig.
 2), and (3) combining both approaches.
-For GTE-Qwen2-7B, the chosen serialization performed best, but all variants outperformed CLMBR-T-Base.
-However, for LLM2Vec-Llama-3.1-8B, only the serialization omitting all values (variant 2) performed slightly better.
-We selected the current serialization for its simplicity and completeness of medical information.
+For GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, the EHR serialization chosen for our experiments performed on par or better than the three alternatives.
+We also selected the current serialization for its simplicity and completeness of medical information.
 This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
-4.4 LLM-Embedding Models and Baselines
-In this study, we evaluated two LLM-embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs.
-These models were selected for their ability to handle the 4.096-token EHR serializations used in our experiments.
+4.4 LLM Embedding Models and Baselines
+In this study, we evaluated two LLM embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs.
+We selected these models for their ability to handle the 4,096-token EHR serializations used in our experiments.
 For comparison, we also tested a smaller variant of both models.
-As additional baselines we included commonly used encoder-only embedding models with smaller input sizes (512 tokens).
+As additional baselines, we included commonly used encoder-only embedding models with smaller input sizes (512 tokens).
 To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation.
-For all LLM-embedding models and smaller language models used in this study, we used the mean pooling of the last layer as the final embedding [26, 44].
+For all LLM embedding and smaller language models, we used mean pooling of the last layer for the final embedding [27, 46].
 The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction.
@@
 GTE-Qwen2-7B
-This LLM-embedding model is based on the Qwen2-7B-Instruct LLM [31] using a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3.584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct.
+This LLM embedding model is based on the Qwen2-7B-Instruct LLM [32], which uses a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3,584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct.
 It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [30].
@@
 Contrastive learning was applied using a mixture of private datasets to enhance embedding performance.
-The model also incorporates instructions tailored for embedding tasks, supporting a context size of up to 32.000 tokens.
+The model also incorporates instructions tailored for embedding tasks and supports a context size of up to 32,000 tokens.
 GTE-Qwen2-1.5B
 A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536.
-It was also trained using the GTE method [29] and supports a context size of up to 32.000 tokens.
+It was also trained using the GTE method [30] and supports a context size of up to 32,000 tokens.
 LLM2Vec-Llama-3.1-8B
-This model is built upon the Llama-3.1-8B-Instruct LLM [32] with a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4.096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised.
+This model is built upon the Llama-3.1-8B-Instruct LLM [33], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised.
 Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [27].
 This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks.
-The finetuning used curated data from the public E5 dataset [45, 46], containing approximately 1.5 million entries.
-The model supports task-specific instructions and accommodates a context size of up to 128.000 tokens.
+The fine-tuning used curated data from the public E5 dataset [47, 48], containing approximately 1.5 million entries.
+The model supports task-specific instructions and supports a context size of up to 128,000 tokens.
 LLM2Vec-Llama-2-1.3B
@@
 DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [50].
-It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and utilizing Gradient-Disentangled Embedding Sharing.
+It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and using Gradient-Disentangled Embedding Sharing.
 We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
 BERT base/large
-BERT is a well-established text embedding model using an encoder-only transformer trained with the masked language modelling objective [15].
+BERT is a well-established text embedding model using an encoder-only Transformer trained with the masked language modeling objective [15].
 We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased.
@@
 BioClinicalBERT
-CLIMBR-T-Base
-CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million deidentified EHRs from Stanford Medicine with autoregressive next code prediction [13, 30].
+CLMBR-T-Base
+CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine with autoregressive next code prediction [13, 31].
 It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base.
 The model has 141M parameters and allows for a context window of 496 codes.
-CLIMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [30].
-It serves as a main baseline for our experiments to test specialized EHR models against general purpose text embeddings models for representing EHR records.
-LLM-Embedding Model and CLIMBR-T-Base
-To test whether the LLM-embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction.
-To this end we simply appended both embeddings.
-The resulting embeddings have dimensions 4.352 (GTE-Qwen2-7B) and 4.864 (LLM2Vec-Llama-3.1-8B).
-Counts-based Model
-Counts models have proven to be strong baselines for EHR prediction tasks [6, 12, 13].
+CLMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [31].
+It serves as a main baseline for our experiments to test specialized EHR models against general-purpose text embedding models for representing EHR records.
+LLM Embedding Model and CLMBR-T-Base
+To test whether the LLM embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction.
+To this end, we simply appended both embeddings.
+The resulting embeddings have dimensions 4,352 (GTE-Qwen2-7B) and 4,864 (LLM2Vec-Llama-3.1-8B).
+Count-based Model
+Count models have proven to be strong baselines for EHR prediction tasks [6, 12, 13].
 The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept.
 We used the counts baseline introduced in [31] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
-Based on the embeddings or the counts vectors given by the methods described above a classification head was trained and validated for each prediction task.
+Based on the embeddings or the counts vectors generated by the methods described above, a classification head was trained and validated for each prediction task.
 For the embedding models we used a logistic regression head.
-For the counts-based model, we used a gradient boosting machine [39], which proved superior [30].
+For the count-based model, we used a GBM [41], which proved superior [31].
 We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
-4.5 Instructions for LLM-Embedding Models
+4.5 Instructions for LLM Embedding Models
 The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts.
 Hence, we added simple instructions for each prediction task based on their respective instruction templates.
-For instance, for prediction of anemia we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”.
+For instance, for prediction of anemia, we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”.
 The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons.
-However, to support task-specific instructions, we had to change this default behavior leading to 1.161.412 instead of 406.379 EHRs to encode resulting in longer processing times.
-The difference between 1.161.412 labels used in our experiments and the total number of labels of 1.178.665 (Table 2) is since some labels even have the same tasks and prediction time and are merged.
-We list all instructions in Table 8 and perform ablations testing the effect of the instructions.
+However, to support task-specific instructions, we had to change this default behavior, leading to 1,161,412 instead of 406,379 EHRs to encode, resulting in longer processing times.
+The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (Table 2) is because some labels share the same task and prediction time and are therefore merged.
+For assignment of new diagnoses tasks in UKB, the same prompt was used for all diseases to reduce computational resources.
+We list all instructions in Table 8 and perform ablations to test the effect of the instructions.
 4.6 Computational Setup and Running Times
 All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems).
-Running the GTE-Qwen2-7B model with our serialization truncated at 4,096-tokens on an 8-GPU DGX system required approximately 20 hours.
+Running the GTE-Qwen2-7B model with our serialization truncated at 4,096 tokens on an 8-GPU DGX system required approximately 20 hours.
 For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset.
@@
 Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs.
-4.7 Performance Results on 15 Prediction Tasks and Few-Shot Setting
+The calculation of the embeddings for the 387,464 UKB patients took approximately 35 hours on a single GPU per task for the LLM-based methods.
+For CLMBR-T-Base, the calculation was faster, only requiring around eight hours.
+4.7 Performance Results on EHRSHOT Prediction Tasks and Few-Shot Setting
 Following the EHRSHOT benchmark, we evaluated all models across 15 prediction tasks under various few-shot settings.
-The benchmark includes a modular pipeline designed to execute key tasks, with the flexibility to optionally utilize a Slurm cluster for distributed execution.
+The benchmark includes a modular pipeline designed to execute key tasks, with the flexibility to optionally use a Slurm cluster for distributed execution.
 Running all steps within this pipeline ensures full reproducibility of results.
-Step four of the pipeline, which generates EHR representations with CLIMBR-T-Base and the counts-based model, was extended to incorporate our method for creating language model-based EHR representations.
+Step four of the pipeline, which generates EHR representations with CLMBR-T-Base and the count-based model, was extended to incorporate our method for creating language model-based EHR representations.
 This adaptation allowed us to reuse significant portions of the existing code, including the task evaluation framework.
 Additionally, we implemented new functionality for EHR serialization and slightly modified other steps of the benchmark to accommodate our experimental setup.
-For instance, the label creation process was adjusted (step three) to enable task-specific instructions for the LLM-embedding models.
+For instance, the label creation process was adjusted (step three) to enable task-specific instructions for the LLM embedding models.
 All modifications have been documented and can be tracked in our public GitHub repository.
-4.8 Effect of Different Contexts Sizes
-We investigated the impact of varying context sizes in the LLM-embedding models.
+4.8 External Validation on UK Biobank
+External validation was performed using data from the UKB, a large-scale prospective cohort study comprising 502,489 UK participants recruited between 2006 and 2010, with a median follow-up of 13.8 years.
+We used linked EHR data from primary care (General Practitioner, GP) and secondary care (Hospital Episode Statistics, HES), providing information on diagnoses, procedures, and prescriptions.
+Initial data preprocessing, including cleaning, feature extraction, missing value imputation, and endpoint selection followed the methodology described in [34].
+All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from the provider and country-specific non-standard vocabularies to OMOP standard vocabularies.
+Participants lacking any recorded GP or HES events either before or after their recruitment date were excluded, resulting in a validation cohort of 387,464 individuals.
+Diagnostic codes were mapped to Phecodes X [53, 54] (derived directly from source ICD-10 or by mapping from SNOMED to ICD-10 codes and subsequently to Phecodes X) primarily for standardized endpoint definition and cohort selection.
+To avoid redundancy with source codes used as features, the Phecodes themselves were excluded during the creation of patient sequences for model input.
+Due to significant challenges in mapping and harmonizing UKB laboratory values [55], and to ensure comparability across models and tasks, laboratory data were excluded entirely, differing from the use of binary labels in [34].
+The final feature set comprised conditions (SNOMED, CVX), medications (RxNorm), and procedures (SNOMED).
+Given that UKB represents a general population cohort, we adjusted the proposed prediction tasks to define relevant longitudinal health trajectory tasks.
+These tasks included: (1) prediction of all-cause hospitalization within the next year (operational outcomes), (2) prediction of incident diagnoses for a set of selected conditions (assignment of new diagnoses), and (3) prediction of all-cause mortality (mortality prediction).
+The selection of incident diseases for the assignment of new diagnoses task group largely followed [34], focusing on common conditions, diseases lacking established risk stratification tools, and specific cardiovascular conditions.
+From the initial 24 endpoints proposed in [34], we treated all-cause death as a separate task.
+For the assignment of new diagnoses and mortality tasks, patients with a diagnosis of the respective endpoint recorded prior to their UKB recruitment date were excluded to prevent data leakage; this exclusion was not applied to the hospitalization task due to high incidence.
+Following task-specific exclusions, the final sample size available for analysis ranged from 300,344 to 359,250 patients, depending on the prediction task (see Table 6 for details).
+For all three task groups, the prediction window was set to one year after the prediction year.
+The pre-trained CLMBR-T-Base model operates with a fixed vocabulary (65,536 total, 26,249 unique codes).
+To use this model, we mapped the 50,702 unique medical codes (SNOMED CT, RxNorm, CVX) present in our processed UKB cohort to the CLMBR-T-Base vocabulary.
+This mapping followed the steps similar to those implemented in the FEMR package101010https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts.
+This process successfully mapped 7,969 unique UKB codes to the CLMBR-T-Base vocabulary (in format ontology/code), which means that the remaining codes were implicitly excluded from the input sequences for this model.
+After creating patient timelines, adding information about birth date, ethnicity, and visits, the data was converted into the MEDS standard [56], from which the embeddings were calculated.
+Additionally, UKB ethnicities were converted to the ethnicity groups used by CLMBR-T-Base.
+For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (mainly the convert_patient function111111https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
+Similar to the approach of Wornow et al.
+[31], we performed an ontology expansion to improve the counts baseline by adding direct parent and grandparent concepts.
+This increased the unique code count from 51,677 to 69,850, aiming to capture broader hierarchical relationships between medical concepts.
+However, the resulting high dimensionality posed memory challenges for a standard counts matrix.
+To mitigate this, we implemented feature selection by removing codes present in less than 50 individuals (0.01% of the cohort), reducing the feature space to 24,509 unique codes.
+To ensure a more comparable baseline to the other models, which implicitly incorporate patient context, we explicitly added normalized age at prediction time and coded sex as additional features to the feature matrix used for the count-based baseline.
+Similar to before, all approaches were evaluated using a five-fold cross-validation.
+Within each fold, hyperparameters were optimized on a dedicated validation set of similar sample size to that of the training partition in that fold.
+For each fold, patients were randomly assigned to train, validation, and test sets.
+The test set for each task within a fold was constructed by selecting all available positive cases (capped to 10,000), and an equal number of negative cases after allocating patients to the train/validation sets.
+To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the train set, even when larger test sets were available.
+4.9 Ablations of EHR Serialization
+To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation experiments.
+First, we assessed the impact of task-specific instructions by replacing them with a generic instruction (Table 8) or removing them entirely to determine their influence on the final embeddings.
+Subsequently, we performed additional ablations by systematically excluding specific components of the serialization.
+These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries.
+Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures.
+Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components.
+This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models use structured medical data for prediction tasks.
+4.10 Effect of Different Time Windows
+To examine the influence of recency on predictive performance, we varied the time window preceding the prediction date used during EHR text serialization.
+We evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B across six intervals: one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history.
+For each window, only events occurring within the specified interval before the prediction time were included.
+Hence, only data from the respective time window contributed to the aggregated information and the visit data in the EHR serialization.
+All other aspects of the serialization, including structure, formatting, and instruction prompts remained unchanged to isolate the effect of the temporal window.
+4.11 Effect of Different Context Sizes
+We investigated the impact of varying context sizes in the LLM embedding models.
 We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings.
-Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1.024, 2.048, and 8.192 tokens.
+Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1,024, 2,048, and 8,192 tokens.
 Input tokens exceeding these thresholds were discarded.
@@
 4.12 Effect of Chunked Contexts
-To further explore whether LLM-embedding models effectively interpret the entirety of their input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs.
-For this, the 4.096-token inputs were divided into smaller chunks of sizes 512, 1.024, and 2.048 tokens101010In our implementation we used these chunks sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models..
+To further explore whether LLM embedding models effectively interpret their entire input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs.
+For this, the 4,096-token inputs were divided into smaller chunks of sizes 512, 1,024, and 2,048 tokens121212In our implementation, we used these chunk sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models..
 Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks.
@@
 This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
-4.10 Ablations of EHR Serialization
-To better understand the contribution of various components in the EHR serialization process to the performance of the LLM-embedding models, we conducted a series of ablation experiments.
-First, we assessed the impact of task-specific instructions by removing them entirely to determine their influence on the final embeddings.
-Subsequently, we performed additional ablations by systematically excluding specific components of the serialization.
-These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries.
-Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures.
-Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components.
-This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models leverage structured medical data for prediction tasks.
+Acknowledgements
+Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 437531118 – SFB 1470 and German Federal Ministry of Education and Research (BMBF) Project-ID 01ZZ2317G.
+Declaration of interest
+The authors declare no competing interests.
+Data availability
+The EHRSHOT data is available through gated access via https://doi.org/10.57761/0gv9-nd83.
+UK Biobank data, including all linked routine health records, are publicly available to bona fide researchers upon application at http://www.ukbiobank.ac.uk/using-the-resource.
+In this study, only primary care data not subject to the Government’s Control of Patient Information (COPI) notice was used (UK Biobank Category 3000).
+Code availability
+All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark.
 Appendix A Appendix
+A.1 Additional Experimental Details
+| Attribute | Total Labels (Positive) |
+| Mortality prediction | |
+| Mortality (Death) | 359245 (802) |
+| Operational Outcomes | |
+| Hospitalization | 359250 (68505) |
+| Assignment of New Diagnoses | |
+| Cardiac arrest | 358963 (87) |
+| Abdominal aortic aneurysm | 358926 (73) |
+| Parkinson’s disease (Primary) | 358729 (86) |
+| Aortic stenosis | 358384 (114) |
+| Rheumatic fever and chronic rheumatic heart diseases | 358064 (154) |
+| Mitral valve insufficiency | 357949 (186) |
+| Endocarditis | 357876 (79) |
+| Suicide ideation and attempt or self harm | 357652 (175) |
+| Cerebral infarction [Ischemic stroke] | 357180 (310) |
+| Pulmonary embolism | 357158 (248) |
+| Atrial fibrillation | 357154 (306) |
+| Heart failure | 356657 (435) |
+| Rheumatoid arthritis | 356029 (316) |
+| Chronic obstructive pulmonary disease [COPD] | 354513 (923) |
+| Psoriasis | 353400 (410) |
+| Chronic kidney disease | 353396 (1001) |
+| Pneumonia | 353232 (802) |
+| Myocardial infarction [Heart attack] | 351970 (619) |
+| Diabetes mellitus | 344185 (1841) |
+| Ischemic heart disease | 341574 (1788) |
+| Anemia | 343235 (1814) |
+| Back pain | 300529 (3933) |
+| Hypertension | 300334 (7653) |
 | Medical Concept | Semantic Codes | Unit | Min-Max Range | Normal Range | Formatting |
@@
 | Task | Prompt |
-| Prefix | Given a patient’s EHR in Markdown format, retrieve relevant passages that answer the query: |
+| Prefix (for all tasks) | Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: |
+| EHRSHOT | |
 | Long Length of Stay | will the patient stay in the hospital for more than 7 days |
@@
 | Lupus | has the patient lupus |
-| Acute MI | Has the patient had an acute myocardial infarction |
+| Acute MI | has the patient an acute myocardial infarction |
 | Chest X-Ray Findings | what are the chest x-ray findings of the patient |
+| Generic (ablation) | what are the key clinical features of the patient to predict future medical events |
+| UK Biobank | |
+| Mortality prediction | will the patient die in the next 1 year |
+| Hospitalization | will the patient be hospitalized in the next 1 year |
+| Assignment of New Diagnoses | has the patient a medical condition |
+A.2 Effect of Different Time Windows
+To assess how the recency of medical history affects prediction performance, we varied the temporal window preceding the prediction time during EHR text serialization.
+The results are summarized in Fig.
+6 (left) and detailed in Table 9.
+GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B exhibited similar overall trends, though GTE-Qwen2-7B performed better with larger time windows.
+The one-week and one-month windows consistently yielded the highest AUROC scores, suggesting that recent clinical data carries the most predictive value for downstream tasks.
+Notably, the 24-hour window produced better results than using the full historical record.
+This trend suggests that recent clinical data often holds the most predictive value, while older information may introduce irrelevant or outdated signals.
+Among the two models, GTE-Qwen2-7B demonstrated stronger performance with longer time windows, indicating a higher capacity to extract relevant features from extended clinical narratives.
+A.3 Effect of Different Context Sizes
+We further tested the impact of varying context sizes on the performance of the LLM embedding models.
+As summarized in Fig.
+6 (right) and Table 10, the two models exhibited distinct behaviors.
+For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4,096 tokens, with only a slight decline observed when using 2,048 or 1,024 tokens.
+This suggests that the most relevant information is contained within the first 1,024 tokens and that the model can accurately extract these details even when presented within a longer context.
+In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2,048 tokens, indicating that the model struggles to effectively handle longer sequences up to 4,096 tokens.
+Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [37].
+Furthermore, extending the context to 8,192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs.
+This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
+A.4 Effect of Chunked Contexts
+To investigate whether the models can process the full 4,096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1,024, and 2,048 tokens.
+For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 11).
+For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4,096-token input remains effectively used even when segmented.
+Notably, using 512-token chunks yielded an overall AUROC performance of 0.731 (0.706-0.756), compared to 0.671 (0.647-0.695) when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints.
+In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes.
+This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2,048 tokens.
+A.5 Detailed Ablation Results for EHRSHOT
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| GTE-Qwen2-7B | |||||
+| Full patient history | |||||
+| 3 years (1,095 days) | |||||
+| 1 year (365 days) | |||||
+| 1 month (30 days) | |||||
+| 1 week (7 days) | |||||
+| 1 day | |||||
+| LLM2Vec-Llama-3.1-8B | |||||
+| Full patient history | |||||
+| 3 years (1,095 days) | |||||
+| 1 year (365 days) | |||||
+| 1 month (30 days) | |||||
+| 1 week (7 days) | |||||
+| 1 day | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| GTE-Qwen2-7B | |||||
+| 8.192 context size | |||||
+| 4.096 context size | |||||
+| 2.048 context size | |||||
+| 1.024 context size | |||||
+| 512 context size | |||||
+| LLM2Vec-Llama-3.1-8B | |||||
+| 8.192 context size | |||||
+| 4.096 context size | |||||
+| 2.048 context size | |||||
+| 1.024 context size | |||||
+| 512 context size | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| GTE-Qwen2-7B | |||||
+| 1 x 4.096 token chunks | |||||
+| 2 x 2.048 tokens chunks | |||||
+| 4 x 1.024 tokens chunks | |||||
+| 8 x 512 tokens chunks | |||||
+| LLM2Vec-Llama-3.1-8B | |||||
+| 1 x 4.096 token chunks | |||||
+| 2 x 2.048 tokens chunks | |||||
+| 4 x 1.024 tokens chunks | |||||
+| 8 x 512 tokens chunks | |||||
+A.6 Additional Performance Results on EHRSHOT
+A.7 Additional Performance Results on UK Biobank

### v4

@@
 Recently, domain-specific EHR foundation models trained on large volumes of unlabeled EHR data have shown improved predictive accuracy and generalization.
-However, their development is constrained by limited access to diverse, high-quality datasets, and by inconsistencies in coding standards and clinical practices.
+However, their development is constrained by limited access to diverse, high-quality datasets, and inconsistencies in coding standards and clinical practices.
 In this study, we explore the use of general-purpose Large Language Models (LLMs) to encode EHR into high-dimensional representations for downstream clinical prediction tasks.
-We convert structured EHR data into markdown-formatted plain text documents by replacing medical codes with natural language descriptions.
+We convert structured EHR data into Markdown-formatted plain-text documents by replacing medical codes with natural language descriptions.
 This enables the use of LLMs and their extensive semantic understanding and generalization capabilities as effective encoders of EHRs without requiring access to private medical training data.
 We show that LLM-based embeddings can often match or even surpass the performance of a specialized EHR foundation model, CLMBR-T-Base, across 15 diverse clinical tasks from the EHRSHOT benchmark.
-To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, a population distinct from that used to train CLMBR-T-Base.
-Notably, one of the tested LLM-based models achieves superior performance for disease onset, hospitalization, and mortality prediction, highlighting robustness to shifts in patient populations.
-Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and generalizable alternative to domain-specific models for clinical prediction.
+Critically, our approach requires no institution-specific training and can incorporate any medical code with a text description, whereas existing EHR foundation models operate on fixed vocabularies and can only process codes seen during pretraining.
+To demonstrate generalizability, we further evaluate the approach on the UK Biobank (UKB) cohort, out-of-domain for CLMBR-T-Base, whose fixed vocabulary covers only 16% of UKB codes.
+Notably, an LLM-based model achieves superior performance for prediction of disease onset, hospitalization, and mortality, indicating robustness to population and coding shifts.
+Our findings suggest that repurposed general-purpose LLMs for EHR encoding provide a scalable and flexible alternative to domain-specific models for clinical prediction.
 keywords:
@@
 EHR data is characterized by variable-length sequences of patient visits, irregular sampling intervals, missing entries, heterogeneous and noisy information, and a wide range of hierarchical medical concepts [11].
-As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based methods [6, 12, 13].
-To mitigate these issues, recent approaches have employed large-scale foundation models that are pre-trained on unlabeled EHR data using unsupervised learning [14].
-Many of these models adopt strategies from natural language processing, such as masked-word prediction as in BERT [15] or autoregressive next-word prediction as in GPT [16].
-Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13].
-However, these techniques also face significant limitations: coding standards and healthcare practices differ strongly across sites, and interoperable EHR foundation models would likely need to be trained on a wide variety of EHR datasets, which is difficult to achieve due to the sensitivity of healthcare data.
-Therefore, the development of EHR-specific foundation models remains constrained by the limited size and restricted availability of EHR data.
-In contrast, LLMs benefit from training on vast general-purpose text corpora and a broad range of natural language tasks [20, 21].
-This extensive pre-training enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications.
-Consequently, LLMs have demonstrated strong performance in extracting medical concepts [22], summarizing medical texts [23], and predicting medical outcomes [24] even in low-resource settings.
-However, most modern LLMs, such as GPT [25] or Llama [26], use a decoder-only transformer architecture, which complicates the generation of robust text representations.
-To overcome this limitation, recent work has introduced methods to convert decoder-only LLMs into effective embedding models for downstream prediction tasks [27, 28, 29, 30].
-Additionally, these state-of-the-art models offer an increased context window, making them well-suited for handling long inputs such as serialized EHR data.
+As a result, deep learning models often achieve only modest improvements over traditional methods such as logistic regression or tree-based models for EHR prediction tasks [6, 12, 13].
+To mitigate these issues, recent approaches have employed large-scale foundation models that are pretrained on unlabeled EHR data using unsupervised learning [14].
+Many of these models adopt strategies from natural language processing, such as masked word prediction as in BERT [15] or autoregressive next word prediction as in GPT [16].
+Treating EHR data as sequences of medical codes enables analogous methods such as masked code prediction [17, 18, 12, 19] or next code prediction [13, 20, 21].
+However, code-based EHR foundation models face two fundamental obstacles to interoperability and generalization: site-specific coding practices and fixed vocabularies learned during pretraining.
+For example, CLMBR-T-Base [13] supports only 26,249 unique codes from its training corpus, and when applied to the UKB with 50,702 unique medical codes, only 7,969 (16%) could be mapped, leaving 84% of codes unseen by the model.
+Achieving interoperability would require pretraining on diverse EHR datasets from many institutions, which is difficult due to the sensitivity of healthcare data.
+This motivates models that operate on natural-language descriptions of clinical codes, which avoid fixed vocabularies and transfer more readily across institutions.
+LLMs benefit from pretraining on vast general-purpose text corpora and a broad range of natural-language tasks [22, 23].
+This extensive pretraining enables their language comprehension and allows them to capture domain-agnostic patterns that can be adapted for healthcare applications.
+Consequently, LLMs have demonstrated strong performance in extracting medical concepts [24], summarizing medical texts [25], and predicting medical outcomes [26], even in low-resource settings.
+Recent work extends LLMs to structured EHR by serializing records into text and either using model generations for prediction [27, 28, 29, 30, 31, 32] or extracting fixed-dimensional embeddings for downstream classifiers [33, 34, 35].
+While both paradigms can be competitive with common baselines, many prior studies only use short context windows and evaluate on private or emergency-department cohorts, limiting longitudinal coverage and external validity [33, 35, 34] (see Section 4.7).
+Moreover, most modern LLMs, such as GPT [36] and Qwen [37, 38], use decoder-only transformer architectures trained with left-to-right objectives, which are not optimized for representation learning.
+To address this limitation, recent work converts decoder-only LLMs into effective LLM embedding models via contrastive learning or related techniques [39, 40, 41, 42, 43].
+Additionally, these state-of-the-art models offer large context windows, making them well-suited for handling long inputs such as serialized EHR data.
 In this study, we evaluate whether general-purpose LLM embedding models can effectively encode EHR data for clinical prediction [44] (see Fig.
 1).
-To this end, we convert structured EHR records into plain text summaries that highlight key patient information at prediction time.
-Using two state-of-the-art LLM embedding models, GTE-Qwen2-7B-Instruct (GTE-Qwen2-7B) [32, 30] and LLM2Vec-Llama-3.1-8B-Instruct (LLM2Vec-Llama-3.1-8B) [27, 33], we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark.
-We focus on the performance in few-shot settings to evaluate the generalization ability of this approach and conduct extensive ablation studies to identify the specific components that drive the LLM ’s effectiveness.
-Finally, we validate our approach on the UKB for predicting mortality, hospitalization and onset of 23 diseases [34] to assess its robustness and compare the generalizability with EHR-specific foundation models.
+To this end, we convert structured EHR records into plain-text summaries that highlight key patient information at prediction time.
+Using a state-of-the-art LLM embedding model, Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43] with a context size of 8,192 tokens, we generate high-dimensional EHR embeddings that serve as inputs to logistic regression classifiers across 15 clinical tasks from the EHRSHOT benchmark.
+We analyze the performance in the few-shot setting, where the downstream logistic regression is trained with only a small number of labeled examples, to evaluate the generalization of this approach, and we conduct extensive ablation studies to identify the components that drive the LLM ’s effectiveness.
+Finally, we validate our approach on the UKB for predicting mortality, hospitalization, and the onset of 23 diseases [45] to assess robustness and compare its generalizability against EHR-specific foundation models.
 2 Results
 2.1 Experimental Setup
-Our primary analysis was conducted using the EHRSHOT database, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023.
+Our primary analysis was conducted using the EHRSHOT benchmark, which contains EHRs from 6,739 adult patients treated at Stanford Health Care and Lucile Packard Children’s Hospital between 1990 and 2023.
 This dataset includes 921,499 visits and over 41.6 million clinical events [44], and serves as a standardized benchmark with 15 clinical prediction tasks organized into 4 task groups, along with predefined splits and publicly available code.
@@
 Due to structural differences between the datasets, not all EHRSHOT tasks were transferable to the UKB (see Section 4.10).
-We focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 chronic diseases [34].
+We thus focused on predicting one-year risk of hospitalization, mortality, and the onset of 23 diseases [45] for the external validation in the UKB.
 The processed UKB subset used in our study consisted of 387,464 patients, approximately 19.5 million visits, and over 72 million clinical events.
 Table 7 provides task descriptions and label distributions.
-To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 4,096 tokens, prioritizing more recent data.
-For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key LOINC concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high).
+To prepare EHR data for use with LLMs, we serialized structured patient records into plain text formatted in Markdown, with a maximum length of 8,192 tokens, prioritizing more recent data.
+For EHRSHOT, serialization began with a reference prediction date to which all event dates were normalized, followed by demographics, recent values for 24 key Logical Observation Identifiers Names and Codes (LOINC) concepts across Body Metrics, Vital Signs, and Lab Results, including units and classification (low/normal/high).
 These were followed by a summary of patient visits and reverse-chronological listings of visit-level events such as conditions, medications, and procedures.
@@
 2, and a detailed explanation appears in Section 4.2.
-For the UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations.
-We evaluated two instruction-tuned LLM embedding models, GTE-Qwen2-7B [32, 30] and LLM2Vec-Llama-3.1-8B [33, 27], using task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 8).
+For UKB, a similar serialization approach was used, with event dates normalized to each patient’s recruitment date and exclusion of detailed measurements due to data limitations.
+In some experiments, we used a 4,096-token limit to accommodate runtime constraints.
+We evaluated three instruction-tuned LLM embedding models, Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43], GTE-Qwen2-7B-Instruct (Qwen2-Emb-7B) [37, 42], and LLM2Vec-Llama-3.1-8B-Instruct (Llama3.1-LLM2Vec-8B) [48, 39], with a primary focus on the first model due to its recency.
+We used task-specific prompts (e.g., “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”, see Table 9).
 Our primary objective was to assess how general-purpose LLM-based embedding models, trained on public text data, perform in encoding EHRs for clinical prediction tasks.
 For comparison, we included CLMBR-T-Base, a 141-million-parameter autoregressive foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine [13, 44].
+Since the EHRSHOT benchmark was derived from the same hospital system, CLMBR-T-Base can be considered an in-domain baseline for this benchmark.
+Notably, the validation and test splits of the benchmark were kept completely separate from the model pretraining (see Fig.
+1 in [44]).
 For each model, we computed patient-level embeddings and trained a logistic regression classifier on the training split, with hyperparameters selected using a validation set.
-As a baseline, we also included a Gradient Boosted Machine (GBM) classifier trained on count-based features, which has previously demonstrated strong performance on EHR tasks [31].
+Following the EHRSHOT protocol, we used the same simple logistic-regression head for all embedding-based representations to assess embedding quality without adding classifier complexity or overfitting risk in the few-shot setting.
+As a baseline, we also trained a Gradient Boosted Machine (GBM) classifier on counts of medical concepts aggregated over predefined intervals before prediction, a method that has previously shown strong performance on EHR tasks [44].
+To test the effect of a GBM head, we additionally trained a GBM on Qwen3-Emb-8B embeddings, which yielded worse results (see Table 11).
 See Section 4.4 for additional details on model architectures.
+2.2 General-Purpose LLM Embeddings Rival Domain-Specific EHR Models
+Using all training and validation examples, Qwen3-Emb-8B performed nearly on par with the in-domain EHR foundation model CLMBR-T-Base on EHRSHOT, with an area under the receiver operating characteristic curve (AUROC) of 0.761 (0.733–0.788) compared to 0.769 (0.746–0.792).
+The LLM embedding model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3).
+Qwen3-Emb-8B matched or exceeded the GBM-based counts baseline, a strong model for EHR prediction tasks [12, 13].
+To assess whether the models encode complementary information, we concatenated the LLM-based and CLMBR-T-Base embeddings and observed a substantial improvement to 0.784 (0.758–0.811).
+We also evaluated the Multiple Embedding Model for EHR (MEME), which encodes different sections of the EHR serialization separately, yielding 0.771 (0.745–0.797) for Qwen3-Emb-8B and 0.734 (0.705–0.764) for MedBERT [34].
+However, MEME required eight times as many model inferences, and the overall performance of the original MedBERT model remained considerably lower than Qwen3-Emb-8B on the full EHR serialization.
+Smaller model variants showed decreased average AUROC: 0.739 (0.713–0.764) for Qwen3-Emb-4B and 0.722 (0.693–0.752) for Qwen3-Emb-0.6B.
+The best encoder-only language models were BERT variants pretrained on medical data.
+Nevertheless, they performed considerably worse than the LLM embedding models, with 0.707 (0.680–0.734) for ClinicalBERT and 0.694 (0.661–0.727) for MedBERT.
+Results for all models are provided in Table 10.
+We plotted parameter counts versus average performance of all language models with different sizes in Fig.
+3, suggesting that the performance of the Qwen3- and Qwen2-based embedding models scales with model size.
+We only observe a slight performance trend for the encoder-only model DeBERTa, while for BERT the performance declines with larger models.
+However, these encoder-only models were applied with chunked 512-token inputs, which may have contributed to this effect (see Section 4.4).
+Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance.
+Nonetheless, the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
 | Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| LLM-Embedding Models | |||||
-| GTE-Qwen2-7B | |||||
-| GTE-Qwen2-1.5B | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| LLM2Vec-Llama 2 1.3B | |||||
-| LLM-Embedding Model + EHR Foundation Model | |||||
-| GTE-Qwen2-7B + CLMBR-T-Base | |||||
-| LLM2Vec-Llama-3.1-8B + CLMBR-T-Base | |||||
 | Baselines [44] | |||||
 | CLMBR-T-Base | |||||
-| Counts-based + GBM | |||||
-| Encoder Language Models | |||||
-| DeBERTaV3 large | |||||
-| DeBERTa V3 base | |||||
-| BERT large | |||||
-| BERT base | |||||
+| Count-based + GBM | |||||
+| LLM Embedding Models | |||||
+| Qwen3-Emb-8B | |||||
+| Qwen3-Emb-4B | |||||
+| Qwen3-Emb-0.6B | |||||
+| LLM Embedding Model + EHR Foundation Model [44] | |||||
+| Qwen3-Emb-8B + CLMBR-T-Base | |||||
+| Multiple Embedding Model for EHR (MEME) [34] with Linear Head | |||||
+| Qwen3-Emb-8B MEME | |||||
+| MedBERT MEME | |||||
+| Encoder Language Models with Chunked Inputs | |||||
 | ClinicalBERT | |||||
+| MedBERT | |||||
 | Model | Mortality prediction | Operational Outcomes (Hospitalization) | Assignment of New Diagnoses | Macro Avg. Across Task Groups |
-| LLM-Embedding Models | ||||
-| GTE-Qwen2-7B | ||||
-| LLM2Vec-Llama-3.1-8B | ||||
 | Baselines [44] | ||||
 | CLMBR-T-Base | ||||
-| Counts-based + GBM | ||||
-2.2 General-Purpose LLM Embeddings Rival Domain-Specific EHR Models
-Using all available training and validation examples, the GTE-Qwen2-7B model outperformed CLMBR-T-Base in two of the four task categories: lab test result prediction and assignment of new diagnoses (see Table 3).
-In contrast, CLMBR-T-Base achieved higher performance on tasks related to operational outcomes and chest X-ray findings.
-LLM2Vec-Llama-3.1-8B embeddings showed slightly lower performance than GTE-Qwen2-7B, exceeding CLMBR-T-Base only in the assignment of new diagnoses task group.
-Both LLM-based models were often on par with or outperformed the GBM-based baseline, which is considered a strong baseline for EHR prediction tasks [12, 13].
-To assess whether the models encode orthogonal information, we concatenated CLMBR-T-Base embeddings with LLM-based embeddings and observed substantial performance improvements for both LLM models.
-The combined embeddings achieved an average area under the receiver operating characteristic curve (AUROC) of 0.786 (0.762–0.811) for GTE-Qwen2-7B, up from 0.755 (0.730–0.780), and 0.773 (0.747–0.799) for LLM2Vec-Llama-3.1-8B, up from 0.734 (0.707–0.762).
-The smaller GTE-Qwen2-1.5B model achieved an average performance of 0.748 (0.722–0.774), only marginally lower than its larger counterpart, suggesting it is a viable lightweight alternative for EHR embeddings.
-In contrast, LLM2Vec-Llama-2-1.3B performed notably worse than the larger version, potentially due to limitations of the older Llama 2 architecture.
-Figure 3 shows a general trend that larger models tend to yield better performance across tasks.
-Among the evaluated models, CLMBR-T-Base remained the most parameter-efficient relative to predictive performance.
-Nonetheless, it is important to note that the LLM-based embedding models were not specifically optimized for structured EHR data, which may explain their performance differences.
-On the UKB dataset, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all three prediction task groups (see Table 4).
-LLM2Vec-Llama-3.1-8B surpassed CLMBR-T-Base only in the task group related to the assignment of new diagnoses.
-The performance drop of CLMBR-T-Base was anticipated, as the UKB represents an out-of-domain dataset with numerous disease codes not included in the model’s original vocabulary.
-Although we applied a careful mapping of UKB Observational Medical Outcomes Partnership (OMOP) concepts to those supported by CLMBR-T-Base to improve compatibility and predictive performance, many concepts remained unmapped due to CLMBR-T-Base’s limited vocabulary.
-This limitation underscores a key challenge in adapting domain-specific models to external datasets, even with considerable manual effort.
-In contrast, the LLM-based models required only a mapping from clinical concepts to their natural language descriptions, which were directly incorporated into the Markdown-based serialization.
-This simplified integration process substantially reduced implementation overhead and emphasized the strong generalization ability of LLM-based encodings across diverse coding systems and healthcare domains.
+| Count-based + GBM | ||||
+| LLM Embedding Model | ||||
+| Qwen3-Emb-8B | ||||
+On the UKB dataset, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base across all three prediction task groups with an overall performance of 0.748 (0.736–0.760) compared to 0.713 (0.698–0.728) (see Table 4).
+The performance drop of CLMBR-T-Base was expected because UKB is out-of-domain and contains many disease codes not included in the model’s pretraining vocabulary.
+Of UKB ’s 50,702 unique medical codes, only 7,969 (16%) mapped to CLMBR’s fixed vocabulary of 26,249 codes.
+Consequently, 84% of codes and their associated clinical signal were effectively invisible to CLMBR-T-Base.
+We applied careful Observational Medical Outcomes Partnership (OMOP) concept mapping to maximize coverage, but fundamental vocabulary constraints remained.
+In contrast, the LLM-based approaches can utilize all 50,702 codes by mapping each to a natural-language description during serialization.
+This ability to incorporate any code with a semantic description, without retraining, is a practical advantage for cross-institution deployment where coding practices vary substantially.
 2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
@@
 5 illustrates the aggregated AUROC across all subtasks within the four task categories for varying numbers of training examples.
-LLM embedding models performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data.
-Notably, GTE-Qwen2-7B consistently outperformed CLMBR-T-Base across all training sizes for lab test result prediction and assignment of new diagnoses.
-The strong results for lab test predictions indicate that the inclusion of the last three lab values in the EHR serialization provides highly predictive features, which GTE-Qwen2-7B effectively encodes.
-In contrast, LLM2Vec-Llama-3.1-8B showed comparatively weaker performance in the few-shot regime.
-It only surpassed CLMBR-T-Base in a limited subset of tasks, primarily those involving the assignment of new diagnoses.
-Nonetheless, both LLM-based models consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining.
+The LLM embedding model performed strongly in the few-shot setting, suggesting effective transfer from general text pretraining to serialized EHR data.
+Notably, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base across all training sizes for lab-test result prediction and assignment of new diagnoses.
+The strong results for lab-test predictions indicate that inclusion of the last three lab values in the EHR serialization provided highly predictive features, which Qwen3-Emb-8B effectively encoded.
+Additionally, Qwen3-Emb-8B consistently outperformed the count-based baseline across most configurations, underscoring the benefits of extensive pretraining.
 Consistent with previous findings [44], the largest performance improvements over the baseline were observed at intermediate training set sizes.
-As the number of labeled data samples increased, the relative advantage of pre-trained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes.
+As the number of labeled data samples increased, the relative advantage of pretrained LLM-based models diminished, highlighting the reduced marginal gains of foundation models in high-data regimes.
 The area under the precision-recall curve (AUPRC) results are shown in Fig.
-7, with task-specific AUROC and AUPRC results provided in Fig.
+9, with task-specific AUROC and AUPRC results provided in Fig.
 10 and Fig.
 11 in the Appendix.
-In contrast to EHRSHOT, the LLM-based methods outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset.
-For the assignment of new diagnoses, which includes 23 tasks, both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed CLMBR-T-Base across all sample sizes (see Fig.
+In contrast to EHRSHOT, the LLM-based method outperformed CLMBR-T-Base in most prediction tasks on the UKB dataset.
+For the assignment of new diagnoses, which includes 23 tasks, Qwen3-Emb-8B outperformed CLMBR-T-Base across all sample sizes larger than two (see Fig.
 12 in the Appendix).
-For hospitalization and mortality prediction, CLMBR-T-Base showed stronger performance in the very low-data regime.
-However, with more samples the performance of the LLM embedding models increased and GTE-Qwen2-7B outperformed CLMBR-T-Base.
-Figure 5 presents the mean AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences.
-On EHRSHOT, GTE-Qwen2-7B generally matched CLMBR-T-Base in performance, while LLM2Vec-Llama-3.1-8B fell between these models and the count-based baseline.
-On the UKB dataset, LLM-based models and CLMBR-T-Base performed comparably with very limited training data of up to four examples.
-With additional training data, GTE-Qwen2-7B surpassed both CLMBR-T-Base and LLM2Vec-Llama-3.1-8B, while the latter closely approached the performance of CLMBR-T-Base.
+For mortality prediction and operational outcomes, CLMBR-T-Base showed stronger performance in the very low-data regime of up to four training samples.
+However, with more samples the performance of the LLM embedding models increased, and Qwen3-Emb-8B outperformed CLMBR-T-Base.
+Figure 5 presents the AUROC across all task groups for both EHRSHOT and UKB, highlighting the observed differences.
+On EHRSHOT, Qwen3-Emb-8B generally matched CLMBR-T-Base in performance and consistently outperformed the count-based baseline.
+On the UKB dataset, the LLM-based model and CLMBR-T-Base performed comparably with very limited training data of up to four examples.
+With additional training data, Qwen3-Emb-8B consistently outperformed CLMBR-T-Base.
 These findings underscore the capability of general-purpose LLM embeddings to rival or exceed the performance of domain-specific models, particularly in settings characterized by domain shift or limited labeled data.
@@
 15 in the Appendix.
+2.4 Effect of the EHR Serialization on LLM-Based Embedding Performance
 | Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
-| generic instructions | |||||
-| no instructions | |||||
-| no demographics | |||||
-| no aggregated | |||||
-| no visits (both) | |||||
-| no conditions | |||||
-| no medications | |||||
-| no procedures | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| generic instructions | |||||
-| no instructions | |||||
-| no demographics | |||||
-| no aggregated | |||||
-| no visits (both) | |||||
-| no conditions | |||||
-| no medications | |||||
-| no procedures |
-2.4 Effect of Serialization Components on LLM-Based EHR Embedding Performance
-To evaluate the contribution of individual serialization components, we conducted ablation studies for both LLM embedding models (see Table 5).
-Replacing task-specific instructions with a generic prompt led to a marked performance decline, reducing average AUROC from 0.755 (0.730-0.780) to 0.739 (0.713-0.765) for GTE-Qwen2-7B and from 0.734 (0.707-0.762) to 0.724 (0.696-0.752) for LLM2Vec-Llama-3.1-8B.
-Removing the instruction entirely caused a further drop for GTE-Qwen2-7B to 0.724 (0.698-0.749), whereas LLM2Vec-Llama-3.1-8B showed minimal change, suggesting greater reliance on instruction prompting by GTE-Qwen2-7B.
-The most significant drop was observed in lab result prediction tasks, highlighting the importance of guidance via task-specific instructions.
-Aggregated measurements (body metrics, vital signs, and lab values) were crucial, especially for lab test prediction, where their removal reduced AUROC from 0.865 (0.858-0.873) to 0.714 (0.701-0.726) for GTE-Qwen2-7B and from 0.778 (0.768-0.789) to 0.706 (0.693-0.719) for LLM2Vec-Llama-3.1-8B.
-This underscores the predictive value of recent lab results and their effective encoding by the models.
-For conditions like thrombocytopenia, hyperkalemia, and hyponatremia, the performance drop was especially pronounced without this information for GTE-Qwen2-7B (see Fig.
+| EHR Serialization Formats | |||||
+| Markdown (ours) | |||||
+| List events | |||||
+| List events with time | |||||
+| JSON | |||||
+| XML | |||||
+| YAML | |||||
+To assess the impact of our hand-crafted EHR serialization, we compared it with a very simple newline-separated event list ordered from most recent to oldest, with and without timestamps.
+Overall performance decreased only slightly from 0.761 (0.733–0.788) to 0.751 (0.728–0.773) (with timestamps) and 0.751 (0.726–0.776) (without timestamps).
+Task-group performance shifted, with gains for operational outcomes, assignment of new diagnoses, and chest X-ray findings, likely reflecting different information prioritization in the serialization.
+Thus, while a hand-crafted format is beneficial, the embedding method is not highly sensitive to very simple data formats.
+Two variants with reversed ordering performed substantially worse (see Table 11).
+Replacing Markdown with JSON, XML, or YAML yielded minimal differences (see Table 5).
+XML showed the largest decrease, which might be due to its greater formatting overhead reducing effective clinical content.
+We further quantified the contribution of individual serialization components with ablations using the Qwen3-Emb-8B model (see Fig.
+6 and Table 12).
+For these runs we used a 4,096-token limit due to computational constraints.
+Replacing task-specific instructions with a generic prompt slightly reduced performance, and removing instructions led to a further decrease.
+The largest performance drop was observed for lab-test prediction, indicating that task-aligned instructions help the model attend to relevant spans.
+Removing individual components from the full serialization generally had modest effects, suggesting redundancy across data sources.
+A notable exception was aggregated lab results, whose removal caused a large drop in lab-test prediction from 0.866 (0.859–0.873) to 0.719 (0.706–0.731), underscoring dependence on recent lab results.
+Interestingly, assignment-of-new-diagnoses performance improved when condition histories were omitted, driven primarily by celiac disease (0.630 to 0.734) and lupus (0.680 to 0.743), suggesting that long lists of past diagnoses can add noise for difficult-to-ascertain or miscoded conditions.
+In serializations consisting of a single component, no component matched the full EHR, underscoring the value of combining modalities.
+The strongest single component was conditions, achieving an AUROC of 0.707 (0.679–0.736), followed by lab results, medications, and procedures.
+For lab-test prediction specifically, a lab-only serialization outperformed the full EHR, likely because unrelated context was reduced.
+Demographics, body metrics, vital signs, and visits yielded much lower overall macro performance (see Table 12).
+Yet, each showed task-specific utility: demographics aided new-diagnosis assignment, body metrics supported operational outcomes, and vital signs and visits were more informative for chest X-ray findings (see Fig.
+6).
+2.5 Effect of Context Length and Temporal Scope on Embedding Effectiveness
+To assess model sensitivity to input length and recency, we performed ablation studies on temporal window and context size using EHRSHOT data (see Fig.
+7 and Table 13).
+The models differed in optimal context sizes, ranging from 8,192 tokens for Qwen3-Emb-8B, 4,096 for Qwen2-Emb-7B, and 2,048 for Llama3.1-LLM2Vec-8B.
+This pattern suggests that newer LLM embedding architectures are better able to handle longer contexts.
+All models exhibited sharp performance drops at 512 tokens, a common constraint in older language models [49].
+All LLM embedding models improved when restricted to more recent data, with the 1-month window outperforming longer histories.
+Qwen3-Emb-8B and Qwen2-Emb-7B even slightly outperformed CLMBR-T-Base with a macro-AUROC of 0.773 (0.746–0.800) and 0.773 (0.746–0.801) (see Table 14) compared to 0.769 (0.746–0.792) for CLMBR-T-Base using the full data (see Table 3).
+Because CLMBR-T-Base is restricted to 496 input codes, it effectively encodes a limited time horizon.
+To disentangle long-context capacity from effects of time windows, we additionally evaluated the count-based baseline across the same temporal windows.
+It followed the same recency trend, indicating an inherent difficulty in performing prediction with historic data in the EHRSHOT benchmark.
+Using data from the day before prediction remained strong across methods, whereas restricting to the last hour led to a pronounced decline.
+2.6 Fine-Tuned LLM Embedding and LLM Decoder Models
+To further contextualize the use of general-purpose LLMs for clinical prediction, we evaluate decoder-style LLMs that generate text, where token probabilities can be used for classification.
+We extend the prompt with an additional instruction to produce Yes and No tokens leveraged for prediction (see Table 9).
+Because of the length of the serialized EHRs, in-context few-shot prompting was infeasible, and the decoder was evaluated only in the zero-shot prompting setting.
+In addition, we fine-tuned both the LLM embedding model and the decoder using Low-Rank Adaptation (LoRA).
+The decoder experiments train the instruction-tuned causal language model Qwen3-8B to generate direct Yes and No predictions for each serialized EHR, while the encoder experiments pair Qwen3-Emb-8B with a supervised classification head (see Section 4.6).
+Both approaches preserve the EHRSHOT task definitions, operate on identical Markdown-based EHR serializations, and employ the same few-shot sampling strategy across all shot sizes and replicates.
+Due to computational constraints, we used a token limit of 4,096 tokens, and for the Anticipating Lab Test Results task group we only used a uniformly random subset of 4,096 test samples.
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnoses | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| Encoder models | |||||
+| Qwen3-Emb-8B | |||||
+| Qwen3-Emb-8B (LoRA) | |||||
+| Decoder models | |||||
+| Qwen3-8B (Zero-Shot) | |||||
+| Qwen3-8B (LoRA) | |||||
+The zero-shot Qwen3-8B decoder achieves a macro-AUROC of 0.651 (0.593–0.707), outperforming all other model variants for up to eight training examples per class (see Fig.
 8).
-Interestingly, removing aggregated values slightly improved new diagnosis prediction for GTE-Qwen2-7B, suggesting a more focused input may benefit some tasks.
-Visit-level information had a strong influence on operational outcome tasks.
-Its removal led to AUROC reductions from 0.771 (0.747-0.795) to 0.752 (0.728-0.776) for GTE-Qwen2-7B and from 0.753 (0.727-0.778) to 0.734 (0.709-0.759) for LLM2Vec-Llama-3.1-8B.
-However, for LLM2Vec-Llama-3.1-8B, removing visit data unexpectedly improved lab result prediction, suggesting possible difficulty in filtering relevant signals from verbose input.
-Removing demographics, conditions, medications, and procedures had only minor effects, indicating these features were either redundant or contributed little predictive signal in isolation.
-While aggregated information and visit data were key drivers for lab and operational outcomes, prediction of new diagnoses and chest X-ray findings was less sensitive to individual serialization components.
-This suggests these tasks rely on broader contextual patterns rather than specific structured elements.
-2.5 Effect of Temporal Scope and Context Length on Embedding Effectiveness
-To assess model sensitivity to input length and recency, we performed ablations on temporal window, context size, and chunking using EHRSHOT data.
-Both LLM embedding models showed improved performance when restricted to more recent data, with the 1-month window outperforming longer histories.
-This suggests that recent clinical information may carry the most predictive signal, while older records may contribute less relevant or potentially distracting details (see Fig.
-6 and Section A.2).
-The models differed in optimal context size: GTE-Qwen2-7B achieved the best performance at the maximum 4,096-token context, whereas LLM2Vec-Llama-3.1-8B performed best at 2,048 tokens (Fig.
-6 and Table 10).
-Both models suffered large drops at 512 tokens, a common constraint in older language models [37].
-To test whether models processed the entire input in context, we evaluated chunked context representations, where fixed-size segments (e.g., 512–2,048 tokens) of a 4,096-token input were embedded separately and averaged (see Table 11).
-GTE-Qwen2-7B showed little degradation from chunking, indicating that simple averaging can be sufficient to preserve performance even under token limits.
-In contrast, LLM2Vec-Llama-3.1-8B benefited from chunking, suggesting it struggles with longer contiguous sequences and prefers shorter inputs.
-These results highlight differences in model-specific sequence processing and suggest practical approaches for adapting to context length constraints.
-Full tables and additional details are available in Section A.3 and Section A.4.
+This pattern reflects architectural differences between the two approaches: the decoder can directly exploit pretrained logits for Yes and No tokens, whereas embedding variants must learn a classification head from minimal supervision.
+At , the fine-tuned decoder surpasses its zero-shot counterpart (0.659 (0.603–0.714)) and continues to improve with additional data (0.707 (0.653–0.758) at , 0.738 (0.684–0.788) at ).
+The encoder-based LLM embedding model exhibits a complementary trade-off.
+Fine-tuning Qwen3-Emb-8B yields steady improvements (0.526 (0.461–0.590) at to 0.716 (0.657–0.770) at ), yet the LoRA-tuned encoder consistently underperforms the frozen-encoder baseline reported in the main results across all shot sizes.
+We expect that larger shot counts or refined hyperparameters could narrow this gap, though computational limits precluded exploration of higher or extensive hyperparameter sweeps.
+At , the fully fine-tuned Qwen3-8B decoder achieves a macro-AUROC of 0.738 (0.684–0.788), marginally exceeding the fine-tuned embedding model (0.727 (0.676–0.775)), but this modest gain comes at substantially higher optimization and inference costs.
 3 Discussion
-Our study shows that general-purpose LLM embedding models, pre-trained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data.
-Models such as GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 31].
-This is notable, as CLMBR-T-Base was trained on data from the same health system as EHRSHOT.
-The strong performance of the LLM-based encoders on the UKB dataset highlights their generalization capabilities.
-GTE-Qwen2-7B consistently outperformed the domain-specific foundation model with at least 16 training examples, while LLM2Vec-Llama-3.1-8B achieved similar results.
-These results are particularly notable given the challenges posed by EHR data, including limited size, heterogeneity, and domain shift.
-Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pre-training enables LLMs to extract clinically meaningful patterns from serialized EHRs.
-Model performance improved with larger model sizes and more advanced architectures, suggesting further gains are likely with future LLM development.
+Our study shows that general-purpose LLM embedding models, pretrained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data.
+The model Qwen3-Emb-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 44].
+This is notable given that CLMBR-T-Base was trained on data from the same health system as EHRSHOT.
+Beyond in-domain performance, the strong results on the UKB dataset highlight the generalization capabilities of LLM-based encoders under population and coding shifts.
+Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pretraining enables LLMs to extract clinically meaningful patterns from serialized EHRs.
+Performance scaled with model size and newer architectures, suggesting further gains are likely with future LLM development.
 In contrast, domain-specific EHR models remain constrained by the scarcity and fragmentation of large, diverse clinical datasets.
-Overall, our findings support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
-Our experiments suggest that several key factors contribute to the success of LLM embedding models for encoding EHR data.
-First, providing clear, task-specific instructions in the serialized EHR text was crucial, guiding models to focus on clinically relevant sections and improving tasks requiring targeted attention.
-This takes advantage of the instruction-tuned nature of LLM embedding models [27, 30], enabling them to extract meaningful patterns from heterogeneous EHR inputs.
-Second, aggregating structured data, particularly recent lab results and visit-level information, substantially improved performance, especially for lab test and operational outcome prediction.
-In contrast, components like demographics, medications, and condition codes contributed less consistently to model accuracy.
-Context size was also critical: GTE-Qwen2-7B effectively handled up to 4,096 tokens, while both models performed poorly with very short or overly long inputs.
-Intermediate time windows, especially those covering one month prior to prediction, yielded optimal performance, suggesting that recent clinical data carries more predictive value than long-term history for many tasks.
-For instance, lab values reflect dynamic physiological states and may lose relevance outside of a narrow temporal window.
-This indicates that LLM embeddings are particularly effective at encoding temporally focused, task-relevant clinical signals.
-Notably, GTE-Qwen2-7B demonstrated greater robustness to extended inputs, highlighting its stronger capacity to extract salient information from longer contexts.
-Finally, combining LLM embeddings with CLMBR-T-Base representations improved overall performance, suggesting that these models capture complementary dimensions of patient history.
-This synergy may stem from CLMBR-T-Base’s limited code-based input window of 496 medical codes [13], or from the broader world knowledge encoded in LLMs that is absent in domain-specific models.
+Overall, these results support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
+Our experiments suggest several key factors contribute to the success of LLM embedding models for encoding EHR data.
+First, task-specific instructions prepended to the EHR serialization improved performance, especially for targets that require focusing on particular spans such as lab results, aligning with the instruction-tuned nature of modern embedding models [39, 42].
+Second, aggregating time series data was critical: removing aggregated laboratory values produced a large drop for lab-test prediction, underscoring the relevance of temporally proximate quantitative measurements.
+At the same time, the approach was robust to the choice of serialization: a very simple newline-separated event list performed only slightly worse than our hand-crafted Markdown serialization.
+Also, alternative data formats such as JSON and YAML showed comparable results, suggesting that model capacity rather than markup specifics drives most of the prediction performance.
+Third, temporal scope and context length mattered.
+Intermediate time windows, for example one month prior to prediction, were consistently optimal across models, indicating that recent history carries disproportionate predictive value.
+Long input contexts of up to 8,192 tokens could only be used effectively by Qwen3-Emb-8B, while older LLM embedding models showed a decrease in performance.
+Finally, we observed practical trade-offs between single-embedding serializations and multi-embedding variants like MEME [34].
+While MEME produced modest average gains for Qwen3-Emb-8B on EHRSHOT, it required roughly eight times more model inferences, raising cost and latency considerations in practice.
 Unlike count-based models and specialized EHR foundation models, LLM-based methods are agnostic to coding systems and data formats.
-Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit their applicability across healthcare settings.
-For instance, CLMBR-T-Base processes only 65,536 codes (26,249 unique), selected based on their entropy from the EHRSHOT dataset [31].
-This restriction may lead to the exclusion of clinically important information when applied to external datasets with unseen codes.
-In contrast, LLMs operate on textual inputs, allowing them to interpret any clinical code mapped to a human-readable description.
-This flexibility is particularly valuable given the ongoing challenges in EHR interoperability due to varying coding practices, privacy constraints, and regulatory barriers, which often hinder the aggregation of large, standardized datasets for model pre-training [38].
-Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [34].
-This mitigates the limitations of count-based models that often exclude infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [39].
-Finally, the ability to represent both structured and unstructured information makes LLMs well-suited for future multimodal EHR applications, including those that integrate clinical notes, diagnostic reports, and imaging metadata [40].
+Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit applicability across healthcare settings.
+For instance, CLMBR-T-Base processes 26,249 unique codes, selected from EHRSHOT [44].
+When applying CLMBR-T-Base to UKB, only 7,969 of 50,702 unique codes, 16%, mapped to this fixed vocabulary, rendering most codes invisible at inference.
+In contrast, LLM encoders operate on text, allowing any clinical code to be incorporated via a human-readable description at serialization time.
+This flexibility directly addresses interoperability challenges arising from site-specific coding practices and privacy constraints that hinder pooled pretraining [50].
+Moreover, broad text pretraining, often including biomedical sources, helps capture the semantics of rare or underrepresented concepts [45], mitigating the bias of count-based approaches that downweight infrequent events [51].
+The same textual interface naturally supports future multimodal extensions that integrate notes, diagnostic reports, and imaging metadata [52].
 Our findings highlight that LLM embedding methods combine the strengths of count-based models, specialized EHR models, and the generalization capabilities of LLMs.
-While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are limited by predefined vocabularies and constrained training data.
-In contrast, LLMs can process any textual representation, enabling them to handle rare or unseen codes, though they often produce unstructured outputs that complicate clinical grounding and calibration [24].
-LLM embedding methods bridge this gap by using the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks.
-By converting EHR data into structured text and embedding it with general-purpose LLMs, we enable the use of standard classifiers such as logistic regression to generate clinically meaningful predictions.
-This approach retains the calibration strengths of traditional models while benefiting from the rich contextual understanding of LLMs.
+Count-based and specialized EHR foundation models can yield well-calibrated probabilities but are constrained by fixed vocabularies and local training distributions.
+Pure generation with LLMs can struggle with grounding and calibration [26].
+Embedding-based pipelines bridge this gap: they leverage the representational power of LLMs while remaining compatible with standard discriminative heads such as logistic regression.
+This approach retains familiar calibration workflows and enables straightforward ensembling with domain-specific representations, as shown by the Qwen3-Emb-8B + CLMBR-T-Base concatenation.
+We compared frozen LLM embeddings, a decoder-style LLM that directly emits Yes and No tokens, and parameter-efficiently fine-tuned variants of both.
+Two key observations emerged.
+First, the zero-shot decoder approach was competitive in the very low-shot regime for up to eight training examples even without task-specific training, likely because pretrained logits for Yes and No tokens can be exploited immediately.
+With more supervision for , LoRA-tuned decoding further improved, reaching a macro-AUROC of 0.738 at .
+Second, end-to-end LoRA fine-tuning of Qwen3-Emb-8B with a linear head produced steady gains with increasing , but did not surpass the frozen-encoder baseline within the explored budgets.
+This supports using a frozen embedding model with a tunable classification head for LLM embedding–based EHR prediction [44].
+Leveraging token outputs from decoder-style LLMs can be a viable alternative when a single, well-defined task justifies the higher training and inference costs, or when strong very-few-shot performance is required but in-context examples are infeasible due to long inputs.
 This study has several limitations.
-First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information, which may introduce bias when comparing against models trained on raw data.
-Additionally, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization.
-While the models achieve strong predictive accuracy across tasks, including few-shot settings, their large parameter counts lead to longer computation time and higher resource demands.
-Moreover, the use of LLM embeddings requires training a separate downstream classifier, foregoing the native zero-shot or few-shot capabilities of LLMs.
-We also limited EHR serialization to 4,096 tokens due to runtime constraints, which may exclude relevant long-term historical data.
-Finally, evaluation was conducted on only two datasets, which may limit the generalizability of our findings across diverse healthcare systems.
-Future research should explore serialization-free approaches that allow LLMs to process raw EHR data directly, reducing biases introduced by manual text transformation.
-Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training.
-Extending the effective context window beyond 4,096 tokens will be important for capturing comprehensive patient histories.
+First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information; while ablations suggest the method is robust to simpler formats, hand-crafted choices can still introduce bias relative to models trained on raw, structured data.
+Second, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization.
+Third, LLM encoders have larger computational footprints than CLMBR-T-Base for population-scale screening, and many analyses were constrained to 4,096–8,192 tokens for runtime reasons, potentially excluding long-term history.
+Fourth, we evaluated two large cohorts, EHRSHOT and UKB; broader external validation, including settings with different documentation practices, laboratory availability, and care pathways, will be necessary to assess generalizability, calibration, and fairness at scale.
+Finally, while our decoder experiments provide a complementary perspective, we did not exhaust the fine-tuning design space, for example data scaling, multi-task objectives, pooling strategies, or calibration tuning, and in-context learning was impractical because serialized EHRs typically filled the context window.
+Future research should explore serialization-free approaches that allow LLMs to process raw EHR tables and timelines, reducing bias from manual text transformation.
+Improved pooling or attention-on-attention mechanisms over long inputs may preserve salient signals without diluting them at extreme context lengths.
+Extending the effective context window beyond 8,192 tokens will be important for capturing comprehensive patient histories.
 Additionally, developing techniques to distill large LLMs into smaller, more efficient models could improve their practical use in clinical settings.
-Expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
+Finally, expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
 4 Methods
@@
 The benchmark uses the Framework for Electronic Medical Records (FEMR) 222GitHub repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data.
-All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark.
-The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [31].
-Task labels are derived from clinical events; hence, a single patient could contribute multiple labels per task, resulting in significant variations in task-specific sample sizes.
-For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses.
-The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pre-trained foundation models [14] due to their ability to generalize from limited training data.
-To this end, the benchmark defines evaluation settings with a constrained number of training and validation examples.
-Specifically, for k in 1, 2, 4, 8, 12, 16, 24, 32, 48, 64, 128, the benchmark uses k positive and k negative training examples, along with k positive and k negative validation examples, to train and tune supervised classifiers.
-Testing is always performed on the full set of examples.
+All extensions and experiments conducted for this paper are publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark.
+The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab-test results, assignment of new diagnoses, and anticipating chest X-ray findings [44].
+Task labels are derived from clinical events, so a single patient can contribute multiple labels per task, resulting in substantial variation in task-specific sample sizes.
+For instance, frequent events like laboratory tests yield many more examples than rarer events such as incident diagnoses.
+The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pretrained foundation models [14].
+Specifically, for , the protocol uses positive and negative training examples and positive and negative validation examples to train and tune supervised classifiers.
+Testing is always performed on the full set of available examples for each task.
 The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and GBMs [53].
@@
 For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [44].
-Macro averages are reported for each task group, and an overall macro average is provided across all groups.
+Mean averages are reported for each task group, and an overall macro average is provided across all groups.
 4.2 EHR Text Serialization
-To use LLM embedding models for representing EHR records, we serialized the records into textual formats.
-We limited the serialization length to 4,096 tokens (approximately 16,000 characters) due to computational constraints (see Section 4.6).
-Two experimental runs were conducted with serializations extending up to 8,192 tokens.
+To use LLM embedding models for representing EHR records, we serialized the records into a textual format agnostic to the prediction task.
+Following the EHRSHOT benchmark, we added all events before the label time.
+We limited the serialization length to 8,192 tokens (approximately 32,000 characters) due to computational constraints (see Section 4.8).
+Some additional experiments were conducted with a 4,096-token limit due to computational constraints.
 The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text.
@@
 We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [57].
-To harmonize dates, all timestamps were normalized relative to January 1, 2024, designated as the prediction reference time.
+To standardize temporal information, all dates were normalized relative to January 1, 2024, the prediction reference time.
+Only past visit dates were included, as they were considered most relevant and helped keep the serialization concise.
+To highlight temporal proximity, each date was annotated with the number of days before the prediction time.
 Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity.
-Since 65% of the dataset comprised time-series data encoded via LOINC, which we found imbalanced, we aggregated LOINC-coded events.
+Since 65% of the dataset comprised time-series data encoded via LOINC, which we found to be imbalanced, we aggregated LOINC-coded events.
 Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts.
@@
 Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable.
-Finally, a detailed chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).
-4.3 Potential Bias of Manually Defining an EHR Text Serialization
-Defining an EHR serialization involved subjective decisions, which may have introduced bias.
-For instance, awareness of the prediction tasks could influence the prioritization of certain data elements, potentially favoring task-relevant information.
-To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records.
-Also, due to computational constraints, we evaluated only three alternatives to our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig.
+Finally, a detailed reverse-chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).
+4.3 Testing Alternative EHR Text Serializations
+Our EHR text serialization contained several design choices and may lack generalizability.
+Specifically, we knew the EHRSHOT prediction tasks before designing the serialization, which could have introduced bias in the design of the EHR serialization.
+In total, we initially evaluated three targeted alternatives with a 4,096-token limit due to computational constraints: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed concepts (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig.
 2), and (3) combining both approaches.
-For GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, the EHR serialization chosen for our experiments performed on par or better than the three alternatives.
-We also selected the current serialization for its simplicity and completeness of medical information.
-This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
+We chose the current serialization due to initial experiments with Qwen2-Emb-7B and Llama3.1-LLM2Vec-8B, where it performed on par or better than the three alternatives.
+To test the effect of our serialization format more broadly, we also evaluated a very simple serialization by creating a newline-separated list of all medical events of a patient ordered by time.
+We tested four variants with different sorting directions and with and without timestamps.
+Additionally, we tested three data formats by converting the structured Markdown serialization into JSON, XML, and YAML.
+We used standard Python libraries for the conversion process.
+Due to different formatting overheads of the formats, the amount of clinical content included under a fixed token budget may differ.
 4.4 LLM Embedding Models and Baselines
-In this study, we evaluated two LLM embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs.
-We selected these models for their ability to handle the 4,096-token EHR serializations used in our experiments.
-For comparison, we also tested a smaller variant of both models.
-As additional baselines, we included commonly used encoder-only embedding models with smaller input sizes (512 tokens).
-To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation.
-For all LLM embedding and smaller language models, we used mean pooling of the last layer for the final embedding [27, 46].
-The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction.
-Below is an overview of all models:
-GTE-Qwen2-7B
-This LLM embedding model is based on the Qwen2-7B-Instruct LLM [32], which uses a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3,584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct.
-It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [30].
-This conversion replaces causal attention with bidirectional attention, enabling the model to attend to both left and right contexts for token embedding, akin to BERT.
+We evaluated three LLM embedding models: Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43], GTE-Qwen2-7B-Instruct (Qwen2-Emb-7B) [37, 42], and LLM2Vec-Llama-3.1-8B-Instruct (Llama3.1-LLM2Vec-8B) [48, 39], based on state-of-the-art decoder-only LLMs.
+We selected these models for their ability to handle the 8,192-token EHR serializations used in our experiments.
+For comparison, we also tested smaller variants of Qwen3-Emb-8B and Qwen2-Emb-7B.
+All LLM-based models received task-specific instructions prepended to the EHR serialization (see Section 4.5), and we used each model’s default embedding configuration.
+For Qwen-based embedding models, we used the last-layer hidden state of the final ([EOS]) token as the patient embedding [37, 43].
+For LLM2Vec models, we applied mean pooling over the last-layer hidden states corresponding to the EHR serialization, excluding the instruction tokens from pooling [39].
+Importantly, the instruction still conditions the internal representation and thus influences the resulting embedding.
+As additional baselines, we included commonly used encoder-only embedding models with shorter input limits (512 tokens).
+To apply them to 8,192-token inputs, we split each input into up to 16 chunks of 512 tokens and averaged the chunk embeddings to obtain a single representation.
+These encoder-only models did not use task instructions, and embeddings were computed via mean pooling over the last hidden layer.
+Below is an overview of all models.
+Qwen3-Embedding-8B/4B/0.6B
+This model family is built on the Qwen3 foundation LLMs [38] and trained as instruction-aware embedding models via a multi-stage recipe [43].
+All variants use decoder-only Transformers with causal attention and produce the embedding from the last-layer hidden state of the final [EOS] token (see Section 4.5).
+The 8B and 4B models have 36 layers, while the 0.6B model has 28 layers; all support a context window of up to 32,000 tokens.
+By default, the embedding dimensionality is 4,096 (8B), 2,560 (4B), and 1,024 (0.6B), with support for flexible output dimensions.
+Training comprises large-scale weakly supervised pretraining on synthetic query–document pairs followed by supervised fine-tuning on labeled and filtered synthetic data optimizing a contrastive learning objective.
+Finally, different checkpoints are merged to improve robustness [43].
+Qwen2-Embedding-7B/1.5B
+Built on GTE-Qwen2-7B-Instruct and GTE-Qwen2-1.5B-Instruct [37], these models use decoder-only Transformers (7B: 28 layers, 28 heads, hidden size 3,584; 1.5B: 28 layers, 12 heads, hidden size 1,536).
+They are first trained with autoregressive next-token prediction and then converted to embedding models via the General Text Embedding (GTE) procedure [42], which replaces the causal mask with a BERT-like bidirectional attention during embedding extraction.
 Contrastive learning was applied using a mixture of private datasets to enhance embedding performance.
-The model also incorporates instructions tailored for embedding tasks and supports a context size of up to 32,000 tokens.
-GTE-Qwen2-1.5B
-A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536.
-It was also trained using the GTE method [30] and supports a context size of up to 32,000 tokens.
-LLM2Vec-Llama-3.1-8B
-This model is built upon the Llama-3.1-8B-Instruct LLM [33], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised.
+The models also incorporate instructions tailored for embedding tasks and support a context size of up to 32,000 tokens.
+Llama3.1-LLM2Vec-8B
+This model is built upon the Llama-3.1-8B-Instruct LLM [48], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096444Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised.
 Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [39].
@@
 The model supports task-specific instructions and supports a context size of up to 128,000 tokens.
-LLM2Vec-Llama-2-1.3B
 DeBERTa v3 base/large
@@
 It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and using Gradient-Disentangled Embedding Sharing.
-We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
+We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively555Hugging Face identifiers: microsoft/deberta-v3-base,large.
 BERT base/large
 BERT is a well-established text embedding model using an encoder-only Transformer trained with the masked language modeling objective [15].
-We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased.
+We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks666Hugging Face identifiers: google-bert/bert-base,large-uncased.
 While not state-of-the-art, BERT models remain widely used in embedding tasks.
 BioClinicalBERT
+MedBERT
+This model builds on BioClinicalBERT via continued domain pretraining on heterogeneous biomedical corpora (N2C2, BioNLP, CRAFT, and biomedical Wikipedia) and was originally introduced for biomedical NER [63].
+We include it as a domain-adapted baseline for comparison888Hugging Face identifier: Charangan/MedBERT.
 CLMBR-T-Base
 CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine with autoregressive next code prediction [13, 44].
-It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base.
+It uses transformers and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base.
 The model has 141M parameters and allows for a context window of 496 codes.
@@
 To this end, we simply appended both embeddings.
-The resulting embeddings have dimensions 4,352 (GTE-Qwen2-7B) and 4,864 (LLM2Vec-Llama-3.1-8B).
+The resulting embeddings have dimensions 4,864 for Qwen3-Emb-8B.
 Count-based Model
@@
 The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept.
-We used the counts baseline introduced in [31] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
+We used the count-based baseline introduced in [44], which uses counts from different time intervals and extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
+Notably, this baseline does not incorporate any values associated with medical concepts, such as laboratory measurements, vital signs, or other continuous fields, but only the occurrence counts of each concept code.
 Based on the embeddings or the counts vectors generated by the methods described above, a classification head was trained and validated for each prediction task.
-For the embedding models we used a logistic regression head.
-For the count-based model, we used a GBM [41], which proved superior [31].
+For the embedding models we used a logistic regression head. For the count-based model, we used a GBM [53], which proved superior [44].
 We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
 4.5 Instructions for LLM Embedding Models
-The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts.
+The Qwen and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts.
 Hence, we added simple instructions for each prediction task based on their respective instruction templates.
@@
 The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons.
-However, to support task-specific instructions, we had to change this default behavior, leading to 1,161,412 instead of 406,379 EHRs to encode, resulting in longer processing times.
-The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (Table 2) is because some labels share the same task and prediction time and are therefore merged.
+However, to support task-specific instructions, we changed this default behavior and encoded each (patient, task, prediction time) instance, leading to 1,161,412 instead of 406,379 EHRs to encode and resulting in longer processing times.
+The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (see Table 2) is because some labels share the same task and prediction time and are therefore merged.
+Encoder-only language models did not use instructions and thus required encoding only the 406,379 EHRs from the original implementation, yielding a substantial computational advantage.
 For assignment of new diagnoses tasks in UKB, the same prompt was used for all diseases to reduce computational resources.
 We list all instructions in Table 9 and perform ablations to test the effect of the instructions.
+4.6 Fine-Tuned LLM Embedding and LLM Decoder Models
+We evaluated two approaches for clinical prediction: LLM embedding models that encode the EHR serialization into a fixed-length vector prior to classification, and LLM decoder models that generate text whose token probabilities can be used for classification.
+We further assessed the effect of fine-tuning for both approaches.
+To this end, we implemented a pipeline for fine-tuning Qwen3-Emb-8B [38, 43] with LoRA adapters [64] applied to the encoder attention projections and MLP blocks.
+The encoder is wrapped in a lightweight classification head consisting of a dropout layer (dropout rate ) and a linear projection.
+Inputs are constructed by concatenating the task instruction and the EHR serialization, tokenized with enforced left padding to preserve relative recency, and batched with a custom collator that truncates to 4,096 tokens.
+The entire model is optimized end-to-end with cross-entropy loss.
+For comparability, both the encoder and decoder variants use the same LoRA hyperparameters (, , dropout ).
+Training uses the Hugging Face Trainer with cosine learning-rate decay, early stopping (patience five epochs), gradient checkpointing, and gradient accumulation chosen so that all variants share an effective batch size of .
+The models are trained with mixed precision (BF16), and zero-shot runs () are evaluated without parameter updates to the backbone.
+The decoder pipeline adapts the instruction-tuned causal Qwen3-8B101010Hugging Face identifier: Qwen/Qwen3-8B checkpoint [38] with LoRA adapters applied to all attention projections and MLP up/down projections (q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj).
+Serialized visits are converted into chat-formatted prompts that additional prepend task instructions enforcing the output of Yes or No tokens (see Table 9).
+During LoRA fine-tuning, the model is trained to predict the literal tokens Yes or No as target labels.
+During scoring, the model predicts token probabilities for Yes and No at the final decoding step We aggregate probabilities across all variants of these tokens (including different capitalizations and punctuation) to yield calibrated decision scores without additional heads.
+For both model families, evaluation computes AUROC, AUPRC, and Brier score on the held-out test set, alongside 1,000 patient-level bootstrap replicates to estimate standard deviations and 95% confidence intervals.
+To keep computational cost manageable, we used a token limit of 4,096 and evaluated the Anticipating Lab Test Results group on a uniformly random subset of 4,096 test samples in each run.
+4.7 Existing Methods Using Language Models for EHR Prediction
+Two broad approaches have emerged for leveraging language models with structured EHR data: using model generations directly for prediction and using vector embeddings derived from the model’s hidden states.
+Early work on the generation-based approach converted structured insurance claims into short text snippets and queried an encoder–decoder model with a 512-token context limit, reporting improvements over logistic regression in few-shot regimes (16–64 examples) for end-of-life, surgery, and length-of-stay tasks [26].
+Subsequent studies explored more capable LLMs, prompt design, and fine-tuning strategies on serialized EHR inputs, generally treating prediction as text generation or calibrated scoring from logits [27, 28, 29, 30, 31, 32].
+One variant constrains the model’s output space to medical codes to better align free-text reasoning with code-based clinical labels, thereby bridging the gap between general-purpose language models and EHR-specific foundation models [65].
+In contrast to generation, embedding-based methods treat an EHR-to-text serialization as input and extract a fixed-dimensional representation for a downstream classifier [33, 34, 35].
+Gao et al.
+provide a systematic comparison across serialization formats, LLMs, and embedding strategies on MIMIC-III and a private 660-patient clinical-deterioration dataset, and report that LLM embeddings paired with a prediction head can be competitive in some settings.
+However, raw numerical features remain strong baselines [33].
+They also evaluate an encoder based approach producing Yes and No tokens in the zero-shot setting but find no discriminatory abilities for Mistral-7B-Instruct and Llama3-8B-Instruct.
+The DeLLiriuM method introduced in [35] also evaluates embeddings derived from different LLMs for delirium prediction and performs detailed model introspection via a SHAP analysis.
+Most similar to our work is the Multiple Embedding Model for EHR (MEME) method, which encodes different EHR modalities such as vitals, medications, and diagnoses into separate embeddings and fuses them with a self-attention layer [34].
+This framework uses the relatively small and efficient MedBERT language model for embedding creation and reports superior performance to GPT-4 [34].
+Our study follows the embedding paradigm but differs from prior work in three ways.
+First, we use LLMs explicitly adapted for embedding via contrastive learning.
+Second, we operate with substantially longer inputs (up to 8,192 tokens), which is important for longitudinal EHR serialization where earlier studies often processed up to 3,076 tokens [33] or relied on chunking [34].
+Third, we evaluate on large public longitudinal cohorts and standardized benchmarks from EHRSHOT and UKB.
+We also include a MEME-style baseline with a linear head in place of self-attention due to data sparsity [34].
+Finally, consistent with [33], we explicitly compare embeddings to direct LLM outputs and examine parameter-efficient fine-tuning to assess robustness and practical trade-offs (see Section 2.6).
 4.8 Computational Setup and Running Times
 All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems).
-Running the GTE-Qwen2-7B model with our serialization truncated at 4,096 tokens on an 8-GPU DGX system required approximately 20 hours.
-For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset.
+Running the Qwen3-Emb-8B model with our serialization truncated at 8,192 tokens on an 8-GPU DGX system required approximately 40 hours.
+For the same setup Qwen2-Emb-7B required approximately 31 hours.
+For the Llama3.1-LLM2Vec-8B model, runtime errors occurred during multi-GPU experiments with the full dataset.
 These issues were resolved by splitting the data into smaller batches, which introduced additional overhead.
-Additionally, we optimized the LLM2Vec code by removing an initial text pruning step that took approximately eight hours when using the full data.
-Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs.
-The calculation of the embeddings for the 387,464 UKB patients took approximately 35 hours on a single GPU per task for the LLM-based methods.
-For CLMBR-T-Base, the calculation was faster, only requiring around eight hours.
+Additionally, we optimized the LLM2Vec code by removing an initial word-boundary token-limit routine that iteratively re-tokenizes to ensure the final cut occurs at a word boundary; this step took approximately eight hours when using the full data.
+With our workaround, inputs are simply truncated at the maximum token number (i.e., the last token may correspond to an incomplete word), which matches how the Qwen implementation handles over-length inputs.
+Importantly, this modification does not perform any semantic text pruning or content selection and should have only a minimal effect on performance, limited to a potentially incomplete trailing word.
+Using this modified setup, the Llama3.1-LLM2Vec-8B model took approximately 46 hours to complete on eight A100 GPUs.
+The calculation of the embeddings for the 387,464 UKB patients took approximately 80 hours on a single GPU per task for Qwen2-Emb-7B and Llama3.1-LLM2Vec-8B and approximately 100 hours for Qwen3-Emb-8B, all for a token length of 8,192.
+For significantly reducing the overall runtime, the calculations can be split onto separate GPUs.
+Additionally, since the embedding for the disease prediction task is not disease specific, for each model only three embeddings had to be calculated per patient.
 4.9 Performance Results on EHRSHOT Prediction Tasks and Few-Shot Setting
@@
 External validation was performed using data from the UKB, a large-scale prospective cohort study comprising 502,489 UK participants recruited between 2006 and 2010, with a median follow-up of 13.8 years.
-We used linked EHR data from primary care (General Practitioner, GP) and secondary care (Hospital Episode Statistics, HES), providing information on diagnoses, procedures, and prescriptions.
-Initial data preprocessing, including cleaning, feature extraction, missing value imputation, and endpoint selection followed the methodology described in [34].
-All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from the provider and country-specific non-standard vocabularies to OMOP standard vocabularies.
+We used linked EHR data from primary care (General Practitioner (GP)) and secondary care (Hospital Episode Statistics (HES)), providing information on diagnoses, procedures, and prescriptions.
+Initial data preprocessing, including cleaning, feature extraction, missing-value imputation, and endpoint selection, followed the methodology described in [45].
+All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from provider and country-specific non-standard vocabularies to OMOP standard vocabularies.
 Participants lacking any recorded GP or HES events either before or after their recruitment date were excluded, resulting in a validation cohort of 387,464 individuals.
@@
 The selection of incident diseases for the assignment of new diagnoses task group largely followed [45], focusing on common conditions, diseases lacking established risk stratification tools, and specific cardiovascular conditions.
-From the initial 24 endpoints proposed in [34], we treated all-cause death as a separate task.
+From the initial 24 endpoints proposed in [45], we treated all-cause mortality as a separate task.
 For the assignment of new diagnoses and mortality tasks, patients with a diagnosis of the respective endpoint recorded prior to their UKB recruitment date were excluded to prevent data leakage; this exclusion was not applied to the hospitalization task due to high incidence.
@@
 For all three task groups, the prediction window was set to one year after the prediction year.
-The pre-trained CLMBR-T-Base model operates with a fixed vocabulary (65,536 total, 26,249 unique codes).
+The pretrained CLMBR-T-Base model operates with a fixed vocabulary of 26,249 unique codes.
 To use this model, we mapped the 50,702 unique medical codes (SNOMED CT, RxNorm, CVX) present in our processed UKB cohort to the CLMBR-T-Base vocabulary.
-This mapping followed the steps similar to those implemented in the FEMR package101010https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts.
+This mapping followed steps similar to those implemented in the FEMR package111111https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts.
 This process successfully mapped 7,969 unique UKB codes to the CLMBR-T-Base vocabulary (in format ontology/code), which means that the remaining codes were implicitly excluded from the input sequences for this model.
@@
 Additionally, UKB ethnicities were converted to the ethnicity groups used by CLMBR-T-Base.
-For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (mainly the convert_patient function111111https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
+For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (primarily the convert_patient function121212https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
 Similar to the approach of Wornow et al.
-[31], we performed an ontology expansion to improve the counts baseline by adding direct parent and grandparent concepts.
+[44], we performed an ontology expansion to improve the count-based baseline by adding direct parent and grandparent concepts.
 This increased the unique code count from 51,677 to 69,850, aiming to capture broader hierarchical relationships between medical concepts.
 However, the resulting high dimensionality posed memory challenges for a standard counts matrix.
-To mitigate this, we implemented feature selection by removing codes present in less than 50 individuals (0.01% of the cohort), reducing the feature space to 24,509 unique codes.
+To mitigate this, we implemented feature selection by removing codes present in fewer than 50 individuals (<0.01% of the cohort), reducing the feature space to 24,509 unique codes.
 To ensure a more comparable baseline to the other models, which implicitly incorporate patient context, we explicitly added normalized age at prediction time and coded sex as additional features to the feature matrix used for the count-based baseline.
-Similar to before, all approaches were evaluated using a five-fold cross-validation.
+All approaches were evaluated using a five-fold cross-validation.
 Within each fold, hyperparameters were optimized on a dedicated validation set of similar sample size to that of the training partition in that fold.
 For each fold, patients were randomly assigned to train, validation, and test sets.
-The test set for each task within a fold was constructed by selecting all available positive cases (capped to 10,000), and an equal number of negative cases after allocating patients to the train/validation sets.
-To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the train set, even when larger test sets were available.
-4.9 Ablations of EHR Serialization
-To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation experiments.
-First, we assessed the impact of task-specific instructions by replacing them with a generic instruction (Table 8) or removing them entirely to determine their influence on the final embeddings.
-Subsequently, we performed additional ablations by systematically excluding specific components of the serialization.
-These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries.
-Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures.
-Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components.
-This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models use structured medical data for prediction tasks.
+The test set for each task within a fold was constructed by selecting all available positive cases (capped at 10,000), and an equal number of negative cases after allocating patients to the train and validation sets.
+To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the training set, even when larger test sets were available.
+4.11 Ablation Studies of EHR Serialization
+To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation studies.
+These ablations used Qwen3-Emb-8B with a 4,096-token limit due to computational constraints.
+We first examined the role of task-specific instructions by replacing them with a generic prompt (see Table 9) or removing instructions entirely, thereby isolating their contribution to the resulting embeddings.
+Next, we systematically excluded individual elements of the serialization, demographics, body metrics, vital signs, laboratory values, visit headings, conditions, medications, and procedures, while keeping all other parts of the pipeline fixed to avoid confounding.
+Finally, we constructed serializations that retained only one of these components at a time to assess its standalone predictive value.
+Together, these experiments identify which elements are most critical for producing effective EHR representations and clarify how the models leverage structured clinical data for downstream prediction tasks.
 4.12 Effect of Different Time Windows
 To examine the influence of recency on predictive performance, we varied the time window preceding the prediction date used during EHR text serialization.
-We evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B across six intervals: one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history.
+We evaluated Qwen3-Emb-8B, Qwen2-Emb-7B, Llama3.1-LLM2Vec-8B, and the count-based baseline with a GBM head across seven intervals: one hour, one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history.
 For each window, only events occurring within the specified interval before the prediction time were included.
@@
 We investigated the impact of varying context sizes in the LLM embedding models.
-We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings.
-Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1,024, 2,048, and 8,192 tokens.
+We sought to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings.
+Specifically, we evaluated Qwen3-Emb-8B, Qwen2-Emb-7B, and Llama3.1-LLM2Vec-8B models with input token limits of 512, 1,024, 2,048, 4,096 and 8,192 tokens.
 Input tokens exceeding these thresholds were discarded.
 Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits.
+All other preprocessing choices and task-specific instructions were held fixed across context-size settings.
 By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
-4.12 Effect of Chunked Contexts
-To further explore whether LLM embedding models effectively interpret their entire input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs.
-For this, the 4,096-token inputs were divided into smaller chunks of sizes 512, 1,024, and 2,048 tokens121212In our implementation, we used these chunk sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models..
-Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks.
-This approach aligns with the mean pooling applied to the last layer that we used to compute embeddings.
-In this setup, task-specific instructions, placed at the beginning of the inputs, were included only in the first chunk.
-This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
 Acknowledgements
@@
 | Generic (ablation) | what are the key clinical features of the patient to predict future medical events |
-| UK Biobank | |
+| UK Biobank (UKB) | |
 | Mortality Prediction | will the patient die in the next 1 year |
@@
 | Assignment of New Diagnoses | has the patient a medical condition |
-A.2 Effect of Different Time Windows
-To assess how the recency of medical history affects prediction performance, we varied the temporal window preceding the prediction time during EHR text serialization.
-The results are summarized in Fig.
-6 (left) and detailed in Table 9.
-GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B exhibited similar overall trends, though GTE-Qwen2-7B performed better with larger time windows.
-The one-week and one-month windows consistently yielded the highest AUROC scores, suggesting that recent clinical data carries the most predictive value for downstream tasks.
-Notably, the 24-hour window produced better results than using the full historical record.
-This trend suggests that recent clinical data often holds the most predictive value, while older information may introduce irrelevant or outdated signals.
-Among the two models, GTE-Qwen2-7B demonstrated stronger performance with longer time windows, indicating a higher capacity to extract relevant features from extended clinical narratives.
-A.3 Effect of Different Context Sizes
-We further tested the impact of varying context sizes on the performance of the LLM embedding models.
-As summarized in Fig.
-6 (right) and Table 10, the two models exhibited distinct behaviors.
-For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4,096 tokens, with only a slight decline observed when using 2,048 or 1,024 tokens.
-This suggests that the most relevant information is contained within the first 1,024 tokens and that the model can accurately extract these details even when presented within a longer context.
-In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2,048 tokens, indicating that the model struggles to effectively handle longer sequences up to 4,096 tokens.
-Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [37].
-Furthermore, extending the context to 8,192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs.
-This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
-A.4 Effect of Chunked Contexts
-To investigate whether the models can process the full 4,096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1,024, and 2,048 tokens.
-For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 11).
-For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4,096-token input remains effectively used even when segmented.
-Notably, using 512-token chunks yielded an overall AUROC performance of 0.731 (0.706-0.756), compared to 0.671 (0.647-0.695) when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints.
-In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes.
-This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2,048 tokens.
-A.5 Detailed Ablation Results for EHRSHOT
+| Additional Prompt Added for Decoder Qwen3-8B | |
+| Decoder Prompt | Answer STRICTLY with a single token: Yes or No. No punctuation, no extra words. |
+A.2 Full Results on EHRSHOT
 | Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
+| Baselines [44] | |||||
+| CLMBR-T-Base | |||||
+| Count-based + GBM | |||||
+| LLM Embedding Models | |||||
+| Qwen3-Emb-8B | |||||
+| Qwen3-Emb-4B | |||||
+| Qwen3-Emb-0.6B | |||||
+| Qwen2-Emb-7B | |||||
+| Qwen2-Emb-1.5B | |||||
+| Llama3.1-LLM2Vec-8B | |||||
+| LLM Embedding Model + EHR Foundation Model [44] | |||||
+| Qwen3-Emb-8B + CLMBR-T-Base | |||||
+| LLM Embedding Model + GBM Head | |||||
+| Qwen3-Emb-8B + GBM | |||||
+| Multiple Embedding Model for EHR (MEME) [34] with Linear Head | |||||
+| Qwen3-Emb-8B MEME | |||||
+| MedBERT MEME | |||||
+| Encoder Language Models with Chunked Inputs | |||||
+| ClinicalBERT | |||||
+| MedBERT | |||||
+| DeBERTaV3 large | |||||
+| DeBERTaV3 base | |||||
+| BERT large | |||||
+| BERT base | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| EHR Serialization Formats | |||||
+| Markdown (ours) | |||||
+| List events | |||||
+| List events old | |||||
+| List events with time | |||||
+| List events old with time | |||||
+| JSON | |||||
+| XML | |||||
+| YAML | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| Original Markdown Serialization with 4,096 Tokens | |||||
+| Full EHR | |||||
+| Instruction Experiments | |||||
+| Generic Instruction | |||||
+| Empty Instruction | |||||
+| Removing Serialization Components | |||||
+| No Demographics | |||||
+| No Body Metrics | |||||
+| No Vital Signs | |||||
+| No Lab Results | |||||
+| No Visits | |||||
+| No Conditions | |||||
+| No Medications | |||||
+| No Procedures | |||||
+| Individual Serialization Components | |||||
+| Only Demographics | |||||
+| Only Body Metrics | |||||
+| Only Vital Signs | |||||
+| Only Lab Results | |||||
+| Only Visits | |||||
+| Only Conditions | |||||
+| Only Medications | |||||
+| Only Procedures | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| Qwen3-Emb-8B | |||||
+| 8,192 context size | |||||
+| 4,096 context size | |||||
+| 2,048 context size | |||||
+| 1,024 context size | |||||
+| 512 context size | |||||
+| Qwen2-Emb-7B | |||||
+| 8,192 context size | |||||
+| 4,096 context size | |||||
+| 2,048 context size | |||||
+| 1,024 context size | |||||
+| 512 context size | |||||
+| Llama3.1-LLM2Vec-8B | |||||
+| 8,192 context size | |||||
+| 4,096 context size | |||||
+| 2,048 context size | |||||
+| 1,024 context size | |||||
+| 512 context size | |||||
+| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
+| Qwen3-Emb-8B (4,096 tokens) | |||||
 | Full patient history | |||||
@@
 | 1 day | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
+| 1 hour | |||||
+| Qwen2-Emb-7B (4,096 tokens) | |||||
 | Full patient history | |||||
@@
 | 1 day | |||||
-| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
-| 8.192 context size | |||||
-| 4.096 context size | |||||
-| 2.048 context size | |||||
-| 1.024 context size | |||||
-| 512 context size | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| 8.192 context size | |||||
-| 4.096 context size | |||||
-| 2.048 context size | |||||
-| 1.024 context size | |||||
-| 512 context size | |||||
-| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnosis | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
-| GTE-Qwen2-7B | |||||
-| 1 x 4.096 token chunks | |||||
-| 2 x 2.048 tokens chunks | |||||
-| 4 x 1.024 tokens chunks | |||||
-| 8 x 512 tokens chunks | |||||
-| LLM2Vec-Llama-3.1-8B | |||||
-| 1 x 4.096 token chunks | |||||
-| 2 x 2.048 tokens chunks | |||||
-| 4 x 1.024 tokens chunks | |||||
-| 8 x 512 tokens chunks | |||||
-A.6 Additional Performance Results on EHRSHOT
-A.7 Additional Performance Results on UK Biobank
+| 1 hour | |||||
+| Llama3.1-LLM2Vec-8B (4,096 tokens) | |||||
+| Full patient history | |||||
+| 3 years (1,095 days) | |||||
+| 1 year (365 days) | |||||
+| 1 month (30 days) | |||||
+| 1 week (7 days) | |||||
+| 1 day | |||||
+| 1 hour | |||||
+| Count-based + GBM | |||||
+| Full patient history | |||||
+| 3 years (1,095 days) | |||||
+| 1 year (365 days) | |||||
+| 1 month (30 days) | |||||
+| 1 week (7 days) | |||||
+| 1 day | |||||
+| 1 hour | |||||
+A.3 Additional Performance Results on EHRSHOT
+A.4 Additional Performance Results on UK Biobank
+| Model | Mortality prediction | Operational Outcomes (Hospitalization) | Assignment of New Diagnoses | Macro Avg. Across Task Groups |
+| Baselines [44] | ||||
+| CLMBR-T-Base | ||||
+| Count-based + GBM | ||||
+| LLM Embedding Models | ||||
+| Qwen3-Emb-8B | ||||
+| Qwen2-Emb-7B | ||||
+| Llama3.1-LLM2Vec-8B | ||||
+A.5 Additional Performance For Encoder and Decoder Models